### v1

# Tony Lindeberg1    Zahra Babaiee2 and Peyman M. Kiasari2

Abstract
This paper presents the results of analysing and modelling a set of 8 “master key filters”, which have been extracted by applying a clustering approach to the receptive fields learned in depthwise-separable deep networks based on the ConvNeXt architecture. For this purpose, we first compute spatial spread measures in terms of weighted mean values and weighted variances of the absolute values of the learned filters, which support the working hypotheses that: (i) the learned filters can be modelled by separable filtering operations over the spatial domain, and that (ii) the spatial offsets of the those learned filters that are non-centered are rather close to half a grid unit. Then, we model the clustered “master key filters” in terms of difference operators applied to a spatial smoothing operation in terms of the discrete analogue of the Gaussian kernel, and demonstrate that the resulting idealized models of the receptive fields show good qualitative similarity to the learned filters.
This modelling is performed in two different ways: (i) using possibly different values of the scale parameters in the coordinate directions for each filter, and (ii) using the same value of the scale parameter in both coordinate directions. Then, we perform the actual model fitting by either (i) requiring spatial spread measures in terms of spatial variances of the absolute values of the receptive fields to be equal, or (ii) minimizing the discrete $l_{1}$- or $l_{2}$-norms between the idealized receptive field models and the learned filters.
Complementary experimental results then demonstrate the idealized models of receptive fields have good predictive properties for replacing the learned filters by idealized filters in depthwise-separable deep networks, thus showing that the learned filters in depthwise-separable deep networks can be well approximated by discrete scale-space filters.
Keywords:
receptive field deep learning discrete continuous Gaussian kernel Gaussian derivative depthwise-separable networks scale space
journal: arXiv preprint
1 Introduction
In computer vision operations, specifically such operations formulated in terms of deep networks applied to image data, a main computational function is performed in terms of the receptive fields, that integrate the image information over non-infinitesimal support regions over the image domain. A both theoretical and practical problem does hence concern how to choose the models to use for such receptive fields in computer vision algorithms.
In the area of scale-space theory, the problem of choosing appropriate models for receptive fields has been addressed from a normative theoretical viewpoint, by formulating assumptions, referred to as scale-space axioms, that reflect desirable properties of a vision system under selected families of symmetry transformations. By axiomatic derivations, it has in this respect been shown that Gaussian kernels and Gaussian derivatives constitute a canonical family of linear filters to be used in the first layer of the image operations, see Iijima (1962), Koenderink (1984), Koenderink and van Doorn (1992), Lindeberg (1996, 2011, 2021a) and Weickert et al. (1999). The idealized models for spatial receptive fields obtained in this way have also been shown to agree qualitatively quite well with biological receptive fields registered by neurophysiological recordings in the retina, the lateral geniculate nucleus (LGN) and the primary visual cortex (V1) of higher mammals, see Lindeberg (2013, 2021a).
In the area of deep learning, the problem of choosing the receptive fields is on the other hand addressed from a pure learning perspective, where the resulting receptive fields then arise as the result of optimizing some suitably selected loss function for a specific deep learning architecture. Hybrid approaches to this problem have also been formulated, where the use of Gaussian derivative operators have been extended to higher layers in deep networks by Jacobsen et al. (2016), Lindeberg (2021b, 2022), Pintea et al. (2021), Sangalli et al. (2022), Penaud-Polge et al. (2022), Gavilima-Pilataxi and Ibarra-Fiallo (2023) and Perzanowski and Lindeberg (2025). The use of Gaussian derivative operators in the higher layers of such hybrid networks is, however, not as theoretically motivated as the use of Gaussian derivatives in the first layer, formally based on axiomatically derived scale-space theory.
Filter 1 Filter 2 Filter 3 Filter 4
Refer to caption Refer to caption Refer to caption Refer to caption
Filter 5 Filter 6 Filter 7 Filter 8
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 1: The original set of 8 “master key filters”, obtained by applying a clustering technique to the receptive fields learned from depthwise-separable deep networks, as extracted by Babaiee et al. (2025a), “through greedy search on the ConvNeXt v2 Tiny model” developed by Liu et al. (2022).
Filter 1 Filter 2 Filter 3 Filter 4
Refer to caption Refer to caption Refer to caption Refer to caption
Filter 5 Filter 6 Filter 7 Filter 8
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 2: Alternative visualization of the original set of 8 “master key filters”, extracted by Babaiee et al. (2025a), while here complemented with a normalization the filters, (i) by rescaling Filters 1–6 to give the same response to matching first-order discrete monomials under discrete convolution as for convolution of the corresponding continuous monomials with corresponding continuous Gaussian derivative operators, according to (63)–(68), and (ii) by adding different constants to Filters 7–8, according to (69)–(73), to minimize the variance-based spatial spread measure of the filters, and finally visualizing the data on a blue-red colour scale with the value 0 corresponding to white, to better reveal the polarities of the filter values.
Idealized 1(A) Idealized 2(A) Idealized 3(A) Idealized 4(A)
Refer to caption Refer to caption Refer to caption Refer to caption
Idealized 5(A) Idealized 6(A) Idealized 7(C2) Idealized 8(A)
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 3: Idealized models of the 8 “master key filters” in Figure 2, in terms of combinations pure difference operators with spatial smoothing filters based on the discrete analogue of the Gaussian kernel, with the amount of spatial smoothing determined from the spatial extents of the 8 “master key filters” using Method A in Section 3.1, based on direct transfer of scale values from the variances for a continuous Gaussian derivative model. (Idealized 1): Non-centered vertical difference operator $(\delta_{y+}T)(m,n;\;\sigma_{x,1},\sigma_{y,1})$. (Idealized 2): Non-centered horizontal difference operator $(\delta_{x-}T)(m,n;\;\sigma_{x,2},\sigma_{y,2})$. (Idealized 3): Non-centered vertical difference operator $(\delta_{y-}T)(m,n;\;\sigma_{x,3},\sigma_{y,3})$. (Idealized 4): Non-centered horizontal difference operator $(\delta_{x+}T)(m,n;\;\sigma_{x,4},\sigma_{y,4})$. (Idealized 5): Centered horizontal difference operator $(\delta_{x}T)(m,n;\;\sigma_{x,5},\sigma_{y,5})$. (Idealized 6): Centered vertical difference operator $(\delta_{y}T)(m,n;\;\sigma_{x,6},\sigma_{y,6})$. (Idealized 7): Sharpening operation $1-\gamma_{7}\,(\nabla_{5}^{2}T)(m,n;\;\sigma_{7},\sigma_{7})$. (Idealized 8): Pure spatial smoothing $T(m,n;\;\sigma_{x,8},\sigma_{y,8})$. (The numerical values of the scale parameters $\sigma_{x,i}$ and $\sigma_{y,i}$ for $i\in\{1,2,3,4,5,6,8\}$ have been determined according to Table 7, and with $\sigma_{7}$ additionally according to Table 13.) (Note that the contrasts of Filters 2 and 3 are reversed in relation to the sign conventions for the corresponding “master key filters”.)
More recently, Babaiee et al. (2024b, 2024a, 2025a) have investigated the properties of depthwise convolutional filters learned across all layers of depthwise-separable convolutional neural networks, as originally proposed by Howard et al. (2017), with particular focus on the ConvNeXt architectures by Liu et al. (2022). Through unsupervised clustering techniques applied to millions of trained filters, they discovered that depthwise convolutional kernels can be categorized into a few distinct classes resembling Gaussian functions and their derivatives. Remarkably, these filter patterns demonstrate generality regardless of layer depth, network architecture, or training data domain. Through systematic greedy search methods, they further demonstrated that this diverse set of learned filters can without major loss of performance be reduced to a smaller set of just 8 so-called “master key filters”, see Figures 12 for two different ways of illustrating the same data.
Notably, these receptive fields do specifically appear to be qualitatively very similar to receptive fields obtained from discrete scale-space theory, in terms of either (see Figure 3 for an illustration):
  • anti-symmetric first-order derivative approximations at moderately fine scales (Filters 5 and 6),
  • a Gaussian blob at a moderately fine scale (Filter 8),
  • a local sharpening operation by subtracting a Laplacian-of-the-Gaussian at a moderately fine scale (Filter 7), or
  • non-centered asymmetric first-order derivative approximations at very fine scales (Filters 1–4).
The subject of this paper is to present a quantitative analysis of these learned filters, and specifically to make comparisons to idealized simplifications with different degrees of simplification in idealized models for spatial receptive fields, based on an earlier presented theory for discrete derivative approximations with scale-space properties in Lindeberg (1993a, 1993b, 2024, 2025).
In summary, the results to be presented will show that the learned filters in depthwise-separable deep networks based on the ConvNeXt architecture can be well modelled by discrete scale-space filters. By quantitative results regarding deep learning, we also show that replacement of the learned filter by idealized scale-space filters in depthwise-separable deep networks also lead to almost as good results as when using the fully learned filters. In these respects, the results presented in this paper generalize the previous solid theoretical foundation of using Gaussian derivative filters as the first layer of linear filters in a vision system to also using (here discrete approximations of) Gaussian derivative operators in the higher layers of deep networks.
Given that the ConvNeXt architecture has been demonstrated to constitute a modern architecture for deep networks with very good ability to compete with Transformer architectures, a main significance of this work is that it establishes a strong link between the filters learned in a reduced ConvNeXt architecture and the idealized filters predicted from theoretically based scale-space theory.
1.1 Structure of this paper
The presentation is organized as follows: Section 2 starts by describing a theoretical background for measuring characteristic properties of the learned filters, based on spatial spread measures defined from spatial means and spatial variances of the absolute values of the filter coefficients and the responses to lower-order polynomials of the learned filters. Section 3 then performs modelling of the 8 “master key filters” using four different types of main methods in terms of:
  • Method A: Spatial spread measures as matched between discrete spatial spread measures of the learned filters and continuous spatial spread measures of continuous Gaussian derivatives.
  • Method B: Spatial spread measures as matched between discrete spatial spread measures of the learned filters and discrete spatial spread measures of discrete scale-space kernels defined by applying small support difference operators to the discrete analogue of the Gaussian kernel.
  • Method C: Minimization of the difference between the learned filters and idealized discrete scale-space filters in discrete $l_{1}$-norm.
  • Method D: Minimization of the difference between the learned filters and idealized discrete scale-space filters in discrete $l_{2}$-norm.
Section 4 first interprets the results obtained in terms of the spaces spanned by the receptive field responses in terms of $N$-jet representations in terms of Gaussian derivatives as well as possible implications for the construction of more general Gaussian derivative networks, with the layers defined from linear combination of Gaussian derivative responses.
Section 5 then applies the fitted idealized discrete scale-space filters as the filtering primitives in depthwise-separable networks applied to the ImageNet dataset, and shows that (i) the idealized filters based on Method B have the predictive properties out of the four main classes of filter modelling methods, and (ii) using the idealized models of the 8 “master key filters” as the computational primitives in depthwise-separable deep networks lead to almost as good results as using the original learned filters. Finally, Section 6 concludes with a summary and discussion of some of the main results.
2 Measuring characteristic properties of the 8 “master key filters” in terms of spatial spread measures and responses to lower-order polynomials
In this section, we will apply a similar methodology, as used for modelling and characterizing the discrete derivative approximation methods in Lindeberg (2024, 2025), to model the 8 “master key filters” extracted by Babaiee et al. (2025a), while also employing structurally closely related methods for determining the parameters of idealized receptive field models, as previously used for matching the parameters between different types of theoretical models for temporal receptive fields in Lindeberg (2023, Sections 3.3–3.4).
2.1 Spatial spread measures for characterizing the spatial extent and the spatial offset of receptive fields
For measuring the overall effect of a non-negative receptive field $h(p)=h(x,y)$ with the image coordinates $p=(x,y)^{T}$ with regard to its spatial extent, it is natural to formulate a spatial spread measure in terms of the spatial mean $M(h(\cdot))$ and the spatial variance $M(h(\cdot,\cdot))$ according to
$\begin{split}M(h(\cdot))&=\frac{\int_{p\in{\mathbb{R}}^{2}}p\,h(p)\,dp}{\int_{p\in{\mathbb{R}}^{2}}h(p)\,dp},\end{split}$ (1)
$\begin{split}V(h(\cdot))&=\frac{\int_{p\in{\mathbb{R}}^{2}}p\,p^{T}\,h(p)\,dp}{\int_{p\in{\mathbb{R}}^{2}}h(p)\,dp}-\left(\frac{\int_{p\in{\mathbb{R}}^{2}}p\,h(p)\,dp}{\int_{p\in{\mathbb{R}}^{2}}h(p)\,dp}\right)^{2},\end{split}$ (2)
over a continuous image domain and according to
$\begin{split}M(h(\cdot))&=\frac{\sum_{p\in{\mathbb{R}}^{2}}p\,h(p)}{\sum_{p\in{\mathbb{R}}^{2}}h(p)},\end{split}$ (3)
$\begin{split}V(h(\cdot))&=\frac{\sum_{p\in{\mathbb{R}}^{2}}p\,p^{T}\,h(p)}{\sum_{p\in{\mathbb{R}}^{2}}h(p)}-\left(\frac{\sum_{p\in{\mathbb{R}}^{2}}p\,h(p)}{\sum_{p\in{\mathbb{R}}^{2}}h(p)}\right)^{2},\end{split}$ (4)
over a discrete image domain.
In previous work, the use of such spatial spread measure has been demonstrated to have a predictive ability to compare the effective amount of spatial smoothing in receptive fields, when performing learning of the scale levels in Gaussian derivative networks using different types of discretization operations for Gaussian derivative operators, see Perzanowski and Lindeberg (2025).
2.1.1 Spatial spread measures applied to zero-order scale-space kernels
For the continuous 2-D Gaussian kernel
$g(x,y;\;\sigma)=g_{\scriptsize\mbox{1D}}(x;\;\sigma)\,g_{\scriptsize\mbox{1D}}(y;\;\sigma)=\frac{1}{2\pi\sigma^{2}}\,e^{-(x^{2}+y^{2})/2\sigma^{2}},$ (5)
with the 1-D Gaussian kernel given by
$g_{\scriptsize\mbox{1D}}(x;\;\sigma)=\frac{1}{\sqrt{2\pi}\sigma}\,e^{-x^{2}/2\sigma^{2}}$ (6)
with standard deviation $\sigma$ and the spatial variance $s=\sigma^{2}$, the resulting variance-based spread measure is then
$V(g(\cdot,\cdot;\;\sigma))=\sigma^{2}\,I,$ (7)
where $I$ denotes the unit matrix. Similarly, for the 2-D discrete analogue of the Gaussian kernel according to (Lindeberg 1990, 2024)
$T(m,n;\;\sigma)=T_{\scriptsize\mbox{1D}}(m;\;\sigma)\,T_{\scriptsize\mbox{1D}}(n;\;\sigma)$ (8)
with the 1-D discrete analogue of the Gaussian kernel given by
$T_{\scriptsize\mbox{1D}}(n;\;\sigma)=e^{-s}I_{n}(\sigma^{2}),$ (9)
where $I_{n}(\sigma^{2})$ denote the modified Bessel functions of integer order (see Abramowitz and Stegun 1964), the corresponding discrete spatial spread measure is correspondingly (Lindeberg 1993a, 2024)
$V(T(\cdot,\cdot;\;\sigma))=\sigma^{2}\,I.$ (10)
Since both the continuous Gaussian kernel and the discrete analogue of the Gaussian kernel are centered, we also have that the spatial mean values are
$\begin{split}M(g(\cdot,\cdot;\;\sigma))=0,\end{split}$ (11)
$\begin{split}M(T(\cdot,\cdot;\;\sigma))=0.\end{split}$ (12)
2.1.2 Spatial spread measures applied to first-order scale-space kernels
Following the methodology in Lindeberg (2024, 2025), we will in this paper also perform corresponding characterizations of receptive fields $h(x,y)$ that are not guaranteed to be non-negative, by then measuring the mean value and the variance of the absolute value of the receptive field $|h(x,y)|$, in terms of the entities $M(|h(\cdot,\cdot)|)$ and $V(|h(\cdot,\cdot)|)$.
By computing these spatial spread measures for the horizontal first-order Gaussian derivative kernel
$g_{x}(x,y;\;\sigma)=\partial_{x}(g(x,y;\;\sigma))=-\frac{x}{\sigma^{2}}\,g(x,y;\;\sigma),$ (13)
we obtain
$\begin{split}M(|g_{x}(\cdot,\cdot;\;\sigma)|)=\left(\begin{array}[]{cc}0\\ 0\end{array}\right),\end{split}$ (14)
$\begin{split}V(|g_{x}(\cdot,\cdot;\;\sigma)|)=\left(\begin{array}[]{cc}2\sigma^{2}&0\\ 0&\sigma^{2}\end{array}\right).\end{split}$ (15)
Similarly, for the vertical first-order Gaussian derivative kernel
$g_{y}(x,y;\;\sigma)=\partial_{y}(g(x,y;\;\sigma))=-\frac{y}{\sigma^{2}}\,g(x,y;\;\sigma),$ (16)
it holds that
$\begin{split}M(|g_{y}(\cdot,\cdot;\;\sigma)|)=\left(\begin{array}[]{cc}0\\ 0\end{array}\right),\end{split}$ (17)
$\begin{split}V(|g_{y}(\cdot,\cdot;\;\sigma)|)=\left(\begin{array}[]{cc}\sigma^{2}&0\\ 0&2\sigma^{2}\end{array}\right).\end{split}$ (18)
Later, we will make use these results in combination with computed discrete variances of discrete kernels, that resemble first-order derivative approximations, for estimating the scale parameter $\sigma$ in corresponding idealized models of receptive fields in terms of scale-space operations.
$\delta_{x+}=\left(\begin{array}[]{ccc}0&0&0\\ 0&-1&+1\\ 0&0&0\end{array}\right)$ $\delta_{x-}=\left(\begin{array}[]{ccc}0&0&0\\ -1&+1&0\\ 0&0&0\end{array}\right)$ $\delta_{y+}=\left(\begin{array}[]{ccc}0&+1&0\\ 0&-1&0\\ 0&0&0\end{array}\right)$ $\delta_{y-}=\left(\begin{array}[]{ccc}0&0&0\\ 0&+1&0\\ 0&-1&0\end{array}\right)$
$\delta_{x}=\left(\begin{array}[]{ccc}0&0&0\\ -1/2&0&+1/2\\ 0&0&0\end{array}\right)$ $\delta_{y}=\left(\begin{array}[]{ccc}0&+1/2&0\\ 0&0&0\\ 0&-1/2&0\end{array}\right)$ $\delta_{xx}=\left(\begin{array}[]{ccc}0&0&0\\ +1&-2&+1\\ 0&0&0\end{array}\right)$ $\delta_{xy}=\left(\begin{array}[]{ccc}-1/4&0&+1/4\\ 0&0&0\\ +1/4&0&-1/4\end{array}\right)$ $\delta_{yy}=\left(\begin{array}[]{ccc}0&1&0\\ 0&-2&0\\ 0&1&0\end{array}\right)$
Figure 4: Computational molecules visualizing the computational functions of the different (top row) non-centered first-order difference operators and (bottom row) centered first-order and second-order difference operators considered in this paper.
2.1.3 Spatial spread measures applied to zero-order pure difference operators
In this paper, we will in the modelling and the analysis stages primarily consider the following non-centered first-order difference operators (see Figure 4 for an illustration):
$\begin{split}(\delta_{x+}h)(m,n)&=h(m+1,n)-h(m,n),\end{split}$ (19)
$\begin{split}(\delta_{x-}h)(m,n)&=h(m,n)-h(m-1,n),\end{split}$ (20)
$\begin{split}(\delta_{y+}h)(m,n)&=h(m,n+1)-h(m,n),\end{split}$ (21)
$\begin{split}(\delta_{y-}h)(m,n)&=h(m,n)-h(m,n-1),\end{split}$ (22)
and the following centered first-order difference operators:
$\begin{split}(\delta_{x}h)(m,n)&=(h(m+1,n)-h(m-1,n))/2,\end{split}$ (23)
$\begin{split}(\delta_{y}h)(m,n)&=(h(m,n+1)-h(m,n-1))/2.\end{split}$ (24)
Additionally, in the following interpretations of the results to be considered in Section 4, we will also consider the following centered second-order difference operators:
$\begin{split}(\delta_{xx}h)(m,n)&=h(m+1,n)-2h(m,n)+h(m-1,n),\end{split}$ (25)
$\begin{split}(\delta_{xy}h)(m,n)&=(h(m+1,n+1)-h(m+1,n-1)\end{split}$
$\begin{split}&-h(m-1,n+1)+h(m-1,n-1))/4,\end{split}$ (26)
$\begin{split}(\delta_{yy}h)(m,n)&=h(m,n+1)-2h(m,n)+h(m,n-1).\end{split}$ (27)
For the non-centered first-order difference operators, we specifically have the following concerning their spatial mean values:
$\begin{split}M(|\delta_{x+}|)=\left(\begin{array}[]{cc}+\tfrac{1}{2}\\ 0\end{array}\right),\end{split}$ (28)
$\begin{split}M(|\delta_{x-}|)=\left(\begin{array}[]{cc}-\tfrac{1}{2}\\ 0\end{array}\right),\end{split}$ (29)
$\begin{split}M(|\delta_{y+}|)=\left(\begin{array}[]{cc}0\\ +\tfrac{1}{2}\end{array}\right),\end{split}$ (30)
$\begin{split}M(|\delta_{y-}|)=\left(\begin{array}[]{cc}0\\ -\tfrac{1}{2}\end{array}\right),\end{split}$ (31)
while for the centered first-order or second-order difference operators, $\delta_{*}\in\{\delta_{x},\delta_{y},\delta_{xx},\delta_{xy},\delta_{yy}\}$, because of their pure mirror symmetries or anti-symmetries, their spatial mean values are all zero:
$M(|\delta_{*}|)=\left(\begin{array}[]{cc}0\\ 0\end{array}\right).$ (32)
For the non-centered first-order difference operators, we specifically have the following concerning their spatial variances:
$\begin{split}V(|\delta_{x+}|)=V(|\delta_{x-}|)=\left(\begin{array}[]{cc}\tfrac{1}{4}&0\\ 0&0\end{array}\right),\end{split}$ (33)
$\begin{split}V(|\delta_{y+}|)=V(|\delta_{y-}|)=\left(\begin{array}[]{cc}0&0\\ 0&\tfrac{1}{4}\end{array}\right),\end{split}$ (34)
the following concerning the centered first-order difference operators:
$\begin{split}V(|\delta_{x}|)=\left(\begin{array}[]{cc}1&0\\ 0&0\end{array}\right),\end{split}$ (35)
$\begin{split}V(|\delta_{y}|)=\left(\begin{array}[]{cc}0&0\\ 0&1\end{array}\right),\end{split}$ (36)
and the following concerning the centered second-order difference operators:
$\begin{split}V(|\delta_{xx}|)=\left(\begin{array}[]{cc}\tfrac{1}{2}&0\\ 0&0\end{array}\right),\end{split}$ (37)
$\begin{split}V(|\delta_{xy}|)=\left(\begin{array}[]{cc}1&0\\ 0&1\end{array}\right),\end{split}$ (38)
$\begin{split}V(|\delta_{yy}|)=\left(\begin{array}[]{cc}0&0\\ 0&\tfrac{1}{2}\end{array}\right).\end{split}$ (39)
Filter $\|h_{i}(\cdot,\cdot)\|_{1}$ $M(|h_{i}(\cdot,\cdot))|$ $V(|h_{i}(\cdot,\cdot))|$
1 2.47 $\left(\begin{array}[]{r}-0.023\\ -0.602\end{array}\right)$ $\left(\begin{array}[]{rr}+1.799&-0.022\\ -0.022&+1.632\end{array}\right)$
2 2.56 $\left(\begin{array}[]{r}+0.559\\ +0.092\end{array}\right)$ $\left(\begin{array}[]{rr}+1.328&+0.012\\ +0.012&+1.425\end{array}\right)$
3 2.73 $\left(\begin{array}[]{r}+0.056\\ +0.693\end{array}\right)$ $\left(\begin{array}[]{rr}+1.584&+0.035\\ +0.035&+1.364\end{array}\right)$
4 2.32 $\left(\begin{array}[]{r}-0.579\\ +0.098\end{array}\right)$ $\left(\begin{array}[]{rr}+1.635&+0.051\\ +0.051&+1.253\end{array}\right)$
5 4.44 $\left(\begin{array}[]{r}+0.057\\ -0.041\end{array}\right)$ $\left(\begin{array}[]{rr}+3.847&-0.011\\ -0.011&+1.848\end{array}\right)$
6 3.81 $\left(\begin{array}[]{r}-0.006\\ +0.086\end{array}\right)$ $\left(\begin{array}[]{rr}+1.634&-0.080\\ -0.080&+3.347\end{array}\right)$
7 1.96 $\left(\begin{array}[]{r}+0.043\\ -0.086\end{array}\right)$ $\left(\begin{array}[]{rr}+1.276&+0.003\\ +0.003&+1.547\end{array}\right)$
8 3.07 $\left(\begin{array}[]{r}-0.029\\ -0.048\end{array}\right)$ $\left(\begin{array}[]{rr}+2.381&-0.020\\ -0.020&+2.356\end{array}\right)$
Table 1: The result of computing the discrete $l_{1}$-norm $\|h(\cdot,\cdot)\|_{1}$ as well as the discrete spatial mean value $M(|h(\cdot,\cdot))|$ according to Equation (3) and the discrete spatial variance $V(|h(\cdot,\cdot))|$ according to Equation (4) for the 8 “master key filters” shown in Figure 1.
2.2 Measuring the spatial extent and the spatial offset of the 8 “master key filters”
Table 1 shows the result of computing the discrete $l_{1}$-norm
$\|h(\cdot,\cdot)\|_{1}=\sum_{(m,n)\in{\mathbb{Z}}^{2}}|h(m,n)|$ (40)
as well as the discrete spatial spread measures $M(|h(\cdot,\cdot))|$ and $V(|h(\cdot,\cdot))|$ according to Equations (3) and (4) for the 8 “master key filters” extracted by Babaiee et al. (2025a) and shown in Figure 1. As can be seen from these results:
  • For all the filters, the off-diagonal elements in the spatial covariance matrices $V(|h(\cdot,\cdot))|$ are much smaller in magnitude than the diagonal elements. In this respect, the learned receptive field appear to be strongly aligned with the Cartesian coordinate directions.
  • For Filters 1–4, one component of the spatial mean value $M(|h(\cdot,\cdot))|$ is very close to zero, while the other component is rather close to either $\tfrac{1}{2}$ or $\tfrac{1}{2}$, thus reflecting a substantial deviation from symmetry. In this respect, there is a close structural similarity to the corresponding spatial mean values for the non-centered first-order difference operators according to Equations (28)–(31).
    There is, however, also a notable difference in the sense that the mean values do not very closely approximate $\pm 0.5$, but rather $\pm 0.6$ or $\pm 0.7$.
  • For Filters 5–6 and Filters 7–8, the spatial mean values $M(|h(\cdot,\cdot))|$ are overall very close to zero, thus indicating a strong degree of either mirror symmetry or antisymmetry.
  • For all the Filters 1–8, the mixed component in the variance-based spatial spread measure $V(|h(\cdot,\cdot))|$ is very close to zero, indicating a very strong connection of the receptive field shapes to the preferred directions of the Cartesian coordinate system.
  • For Filters 5–6, the ratio between the diagonal elements in the variance-based spatial spread measure $V(|h(\cdot,\cdot))|$ is very close to either 2 or $\tfrac{1}{2}$, indicating a rather close similarity to corresponding continuous variance-based spatial spread measures for first-order Gaussian derivative kernels according to Equations (15) and (18). The similarity is, however, not fully perfect, thus reflecting a minor notable difference.
  • For Filters 1–4, the ratio between the diagonal elements in the variance-based spatial spread measure $V(|h(\cdot,\cdot))|$ is, however, not in any way close to either 2 or $\tfrac{1}{2}$, indicating that a direct transfer of the corresponding expressions (15) and (18) for continuous first-order Gaussian derivatives does not appear as straightforward.
  • For Filter 8, the diagonal components of the variance-based spatial spread measure $V(|h(\cdot,\cdot))|$ are very similar, thus with a very good qualitative agreement with Equations (7) and (10).
  • For Filter 7, the diagonal elements of the variance-based spatial spread measure $V(|h(\cdot,\cdot))|$ are somewhat similar, although differing in a clearly notable respect.
Concerning the interpretation of in particular the variance-based spatial spread measures for Filters 7-8, it should, however, be noted that the definitions of these spatial spread measures assume a background magnitude value equal to zero, while Filters 7–8 in the set of “master key filters” have background magnitude levels that are significantly different from zero, thereby leading to substantial bias in the spatial spread measures, by substantially overestimating the amount of spatial extent of the filters. Thus, some modification is needed, to be able to apply the variance-based spatial spread measures $V(|h(\cdot,\cdot))|$ for estimating filter parameters in idealized models to be used for modelling Filters 7–8.
2.3 Measuring the accuracy in relation to centered or non-centered difference operators combined with spatial smoothing
A natural consistency requirement when constructing a centered discrete derivative approximation operator $\Delta_{x^{\alpha}y^{\beta}}$ that is to approximate the derivative operator $\partial_{x^{\alpha}y^{\beta}}$ of a given order $(m,n)\in{\mathbb{Z}}_{+}^{2}$ over a 2-D spatial domain is to require that the result of applying the discrete derivative approximation filter of order $(m,n)$ applied to a monomial of the same order leads to a response that is spatially constant, with specifically
$\Delta_{x^{\alpha}y^{\beta}}(x^{\alpha}\,y^{\beta})=\alpha!\,\beta!,$ (41)
where $\alpha!=\alpha\cdot(\alpha-1)\dots 2\cdot 1$ denotes the factorial of the horizontal derivative order $\alpha$ and $\beta!=\beta\cdot(\beta-1)\dots 2\cdot 1$ denotes the factorial of the vertical derivative order $\beta$. Furthermore, for any integer monomial of order $(a,b)\in{\mathbb{Z}}_{+}^{2}$, if either $a<\alpha$ or $b<\beta$ (or both), we should require that
$\Delta_{x^{\alpha}y^{\beta}}(x^{a}\,y^{b})=0.$ (42)
These consistency requirements are similar to the results of applying corresponding continuous pure derivative operations $\partial_{x^{\alpha}y^{\beta}}$ to corresponding monomials.
2.3.1 Shift-adjusted consistency requirement for non-centered difference operators designed to approximate spatial derivative operators
With regard to modelling the 8 “master key filters” in Figure 1, where we have according to the analysis in connection with Table 1 noted that the 8 “master key filters” (i) are not necessarily centered, and (ii) do not decrease as rapidly with increasing distance from the center as the discrete derivative approximation filters considered in Lindeberg (2024, 2025),  we do for the purpose of the forthcoming analysis in this paper generalize the criterion in Equation (41) into the shift-adjusted consistency requirement
$\Delta_{x^{\alpha}y^{\beta}}((x-m_{x})^{\alpha}\,(y-m_{y})^{\beta})=\alpha!\,\beta!,$ (43)
and generalize the criterion in Equation (42) into the shift-adjusted consistency requirement
$\Delta_{x^{\alpha}y^{\beta}}((x-m_{x})^{a}\,(y-m_{y})^{b})=0,$ (44)
to hold for any $(a,b)\in{\mathbb{Z}}_{+}^{2}$, if either $a<\alpha$ or $b<\beta$ (or both), where $m_{x}$ and $m_{y}$ denote the components in the spatial mean vector of the absolute value of the discrete filter $h_{x^{\alpha}y^{\beta}}(m,n)$, that implements the computational function of the discrete derivative approximation operator $\Delta_{x^{\alpha}y^{\beta}}$ according to
$M(|h_{x^{\alpha}y^{\beta}}(\cdot,\cdot)|)=\left(\begin{array}[]{cc}m_{x}\\ m_{y}\end{array}\right).$ (45)
In the following analysis to be performed, we will specifically evaluate these criteria at the centers of the responses of the discrete derivative approximation filters to different types of polynomials of the form $(x-m_{x})^{a}\,(y-m_{y})^{b}$.
Specifically, these consistency requirements hold for all of the pure non-central or central difference operators $\delta_{x+}$, $\delta_{x-}$, $\delta_{y+}$, $\delta_{y-}$, $\delta_{x}$, $\delta_{y}$, $\delta_{xx}$, $\delta_{xy}$ and $\delta_{yy}$ that we consider in this work.
2.3.2 Responses to monomials for continuous Gaussian derivative operators
Due to the fact that the spatial scale-space representation of any image $f(x,y)$, obtained by Gaussian smoothing expressed in terms of a parameterization of the scale parameter in terms of the variance $s=\sigma^{2}$ of the Gaussian kernel according to
$L(\cdot,\cdot;\;s)=g(\cdot,\cdot;\;s)*f(\cdot,\cdot),$ (46)
satisfies the diffusion equation
$\partial_{s}L=\frac{1}{2}\,\nabla^{2}L=\frac{1}{2}\,(L_{xx}+L_{yy}),$ (47)
if follows that evolution over scale for monomial input $f_{ab}(x,y)=x^{a}\,y^{b}$ can be described in terms of diffusion polynomials, where we for the lowest orders $a$ and $b$ have that
$\begin{split}q_{00}(x,y;\;s)=1,\end{split}$ (48)
$\begin{split}q_{10}(x,y;\;s)=x,\end{split}$ (49)
$\begin{split}q_{01}(x,y;\;s)=y,\end{split}$ (50)
$\begin{split}q_{20}(x,y;\;s)=x^{2}+s,\end{split}$ (51)
$\begin{split}q_{11}(x,y;\;s)=x\,y\end{split}$ (52)
$\begin{split}q_{02}(x,y;\;s)=y^{2}+s.\end{split}$ (53)
From this summary, we can thus read off that the Gaussian smoothing function will leave the constant input function $f_{00}(x,y)=1$ as well as the linear input functions $f_{10}(x,y)=x$ and $f_{01}(x,y)=y$ unchanged.
This property will then imply that the responses of the zero-order Gaussian kernel and its first order derivatives in the $x$- and $y$-directions are given by
$\begin{split}g(\cdot,\cdot)*1&=1,\end{split}$ (54)
$\begin{split}g(\cdot,\cdot)*x&=x,\end{split}$ (55)
$\begin{split}g(\cdot,\cdot)*y&=y,\end{split}$ (56)
$\begin{split}g_{x}(\cdot,\cdot)*1&=0,\end{split}$ (57)
$\begin{split}g_{x}(\cdot,\cdot)*x&=1,\end{split}$ (58)
$\begin{split}g_{x}(\cdot,\cdot)*y&=0,\end{split}$ (59)
$\begin{split}g_{y}(\cdot,\cdot)*1&=0,\end{split}$ (60)
$\begin{split}g_{y}(\cdot,\cdot)*x&=0,\end{split}$ (61)
$\begin{split}g_{y}(\cdot,\cdot)*y&=0.\end{split}$ (62)
By superposition, this property will then also apply to any shift-adjusted input polynomial of the form $f_{ab}(x,y)=(x-m_{x})^{a}\,(y-m_{y})^{b}$.
By further differentiating these relationsship, it follows that the first-order non-centered Gaussian derivative response in the $x$-direction to the monomial input $f_{10}(x,y)=x-m_{x}$ as well as the first-order non-centered Gaussian derivative response in the $x$-direction to the monomial input $f_{01}(x,y)=y-m_{y}$ will be equal to 1, if the centering for the non-centered Gaussian derivative responses is around the point $(m_{x},m_{y})^{T}$.
Filter $h_{i}(\cdot,\cdot)*1$ $h_{i}(\cdot,\cdot)*(x-m_{x})$ $h_{i}(\cdot,\cdot)*(y-m_{y})$
1 $2.0\cdot 10^{-6} -$0.079 +1.556
2 $5.0\cdot 10^{-6} -$1.380 +0.396
3 $2.0\cdot 10^{-6}$ +0.129 -1.091
4 $5.6\cdot 10^{-18}$ +0.839 +0.029
5 $2.0\cdot 10^{-6}$ +7.750 -0.063
6 $4.0\cdot 10^{-6} -$0.160 +6.137
7 $1.0\cdot 10^{-6} +$0.085 -0.169
8 $4.0\cdot 10^{-16} -$0.072 -0.167
Table 2: The result of computing the responses to the monomials $\theta_{00}=1$, $\theta_{10}=x-m_{x}$ and $\theta_{01}=y-m_{y}$ for the 8 “master key filters” shown in Figure 1, where the spatial offset $(m_{x},m_{y})^{T}=M(|h(\cdot,\cdot)|$ represents the spatial mean of the absolute value of each filter.
Filter $h_{i,\scriptsize\mbox{norm}}(\cdot,\cdot)*(x-m_{x})$ $h_{i,\scriptsize\mbox{norm}}(\cdot,\cdot)*(y-m_{y})$
1 -0.051 +1.000
2 +1.000 -0.287
3 -0.118 +1.000
4 +1.000 +0.035
5 +1.000 -0.008
6 -0.026 +1.000
Table 3: The result of computing the responses to the monomials $\theta_{00}=1$, $\theta_{10}=x-m_{x}$ and $\theta_{01}=y-m_{y}$ for normalized versions of the first 6 “master key filters” shown in Figure 1, where the spatial offset $(m_{x},m_{y})^{T}=M(|h(\cdot,\cdot)|$ represents the spatial mean of the absolute value of each filter. For this table, the amplitudes of the original first 6 “master key filters” have been rescaled, to make the response equal to 1 for the monomial of $\theta_{10}=x-m_{x}$ or $\theta_{01}=y-m_{y}$ that best matches an idealized approximation of the filter as either a derivative in the $x$-direction or the $y$-direction.
2.3.3 Responses to monomials for the 8 “master key filters” operators
Table 2 shows the result of computing the responses to the monomials $\theta_{00}=1$, $\theta_{10}=x-m_{x}$ and $\theta_{01}=y-m_{y}$ for the 8 “master key filters” in Figure 1, with the spatial offset $(m_{x},m_{y})^{T}$ determined from the spatial mean of the absolute value of each filter according to $(m_{x},m_{y})^{T}=M(|h(\cdot,\cdot)|$. As can be seen from these results:
  • The responses of the filters to the constant function $\theta_{00}=1$ is very close to zero for all the filters, indicating that the filter are in this respect DC-balanced.
  • For Filters 2, 4 and 5, the largest response is obtained for the monomial $\theta_{10}=x-m_{x}$, indicating that these filters respond as strongest in a way rather closely related to first-order derivatives in the horizontal $x$-direction.
  • For Filters 1, 2 and 6, the largest response is obtained for the monomial $\theta_{01}=y-m_{y}$, indicating that these filters respond as strongest in a way rather closely related to first-order derivatives in the vertical $y$-direction.
  • The Filters 1–6, which approximate first-order derivatives in either the horizontal $x$-direction or the vertical $y$-direction, are, however, not normalized to return derivative responses to the functions $\theta_{10}=x-m_{x}$ and $\theta_{01}=y-m_{y}$, to be equal to one in the respective cases of matching orders between the monomials and the derivative approximations.
  • For Filters 7 and 8, which have the qualitatively similar visual appearance as a discrete Gaussian kernel, or a discrete delta function with a subtracted Laplacian-of-the-Gaussian, the DC-balancing of the “master key filters” is, however, not fully consistent with discrete Gaussian models of those filters, since the both the continuous integral of the continuous Gaussian kernel and the discrete sum of the filter coefficients in discrete approximations of Gaussian kernels should be equal to 1
Table 3 shows the result of an extension of this analysis, where we have for the first 6 “master key filters” normalized the amplitudes of the receptive fields to make the response equal to 1 for the monomial of $\theta_{10}=x-m_{x}$ or $\theta_{01}=y-m_{y}$ that best matches an idealized approximation of the filter as either a derivative in the $x$-direction or the $y$-direction:
$\begin{split}h_{1,\scriptsize\mbox{norm}}(m,n)=\frac{h_{1}(m,n)}{h_{1}(\cdot,\cdot)*(y-m_{y})|_{(x,y)^{T}=(0,0)^{T}}},\end{split}$ (63)
$\begin{split}h_{2,\scriptsize\mbox{norm}}(m,n)=\frac{h_{2}(m,n)}{h_{2}(\cdot,\cdot)*(x-m_{x})|_{(x,y)^{T}=(0,0)^{T}}},\end{split}$ (64)
$\begin{split}h_{3,\scriptsize\mbox{norm}}(m,n)=\frac{h_{3}(m,n)}{h_{3}(\cdot,\cdot)*(y-m_{y})|_{(x,y)^{T}=(0,0)^{T}}},\end{split}$ (65)
$\begin{split}h_{4,\scriptsize\mbox{norm}}(m,n)=\frac{h_{4}(m,n)}{h_{4}(\cdot,\cdot)*(x-m_{x})|_{(x,y)^{T}=(0,0)^{T}}},\end{split}$ (66)
$\begin{split}h_{5,\scriptsize\mbox{norm}}(m,n)=\frac{h_{5}(m,n)}{h_{5}(\cdot,\cdot)*(x-m_{x})|_{(x,y)^{T}=(0,0)^{T}}},\end{split}$ (67)
$\begin{split}h_{6,\scriptsize\mbox{norm}}(m,n)=\frac{h_{6}(m,n)}{h_{6}(\cdot,\cdot)*(y-m_{y})|_{(x,y)^{T}=(0,0)^{T}}}.\end{split}$ (68)
As can be seen from these results:
  • The complementary response to a first-order monomial in the direction orthogonal to the preferred direction of the receptive field is comparably rather low (of the order of $5~\%$ or below) for Filters 1, 4, 5 and 6.
    In this respect, the responses to first-order monomials is rather well consistent as approximations to first-order derivative operators for Filters 1, 4, 5 and 6.
  • For Filter 2, the response to a first-order monomial in the direction orthogonal to the preferred direction of the receptive field is notable (of the order of $12~\%$) for Filter 3, and very notable (of the order of $30~\%$) for Filter 2.
    In this respect, there are notable deviations from approximations to first-order derivative operators for Filter 3 and specifically for Filter 2.
To conclude, Filters 1–6 show qualitatively rather good or very good structural similarities to first-order derivative approximation filters along the Cartesian coordinate direction.
The matches to corresponding fully idealized models in terms of Gaussian derivatives or discrete approximations thereof is, however, far from perfect. A possible source to deviations from the idealized theory could also be because of from variations in the background periphery of the learned filters, which, by a lack of symmetry in those spurious variations, could contribute significantly to the responses to the first-order monomials.
2.4 Reducing the bias in the estimates of the spatial variance of the Gaussian-smoothing-like receptive fields by DC-compensation
As previously remarked, for Filters 7 and 8, the estimates of the spatial extents of the filters will be significantly biased, because of the lack of a determination of the DC-level of these filters to correspond to a background level of about zero. Instead, as we saw from the experimental results in Table 2, the response of these filters to the constant function $\theta_{00}=1$ is very close to zero, as opposed to being close to one, as it should be for a continuous Gaussian filter or a discrete approximation thereof.
To address this issue, we will in the following DC-correct Filters 7 and 8 according to
$\begin{split}h_{7,\scriptsize\mbox{DC}}(m,n;\;C_{7})=h_{7}(m,n)-C_{7},\end{split}$ (69)
$\begin{split}h_{8,\scriptsize\mbox{DC}}(m,n;\;C_{8})=h_{8}(m,n)-C_{8},\end{split}$ (70)
by determining the values of the DC-correction-constants $\hat{C}_{7}$ and $\hat{C}_{8}$ so as to minimize the determinant of the variance-based spatial spread measure $V(|h(\cdot,\cdot)|)$ according to
$\hat{C_{i}}=\operatorname{argmin}_{C_{i}}\det V(|h_{i}(\cdot,\cdot)-C_{i}|),$ (71)
for $i\in\{7,8\}$, and then further normalize the resulting DC-corrected filters according to
$\begin{split}h_{7,\scriptsize\mbox{norm}}(m,n)=\frac{h_{7,\scriptsize\mbox{DC}}(m,n;\;\hat{C}_{7})}{h_{7,\scriptsize\mbox{DC}}(\cdot,\cdot;\;\hat{C}_{7})*1|_{(x,y)^{T}=(0,0)^{T}}},\end{split}$ (72)
$\begin{split}h_{8,\scriptsize\mbox{norm}}(m,n)=\frac{h_{8,\scriptsize\mbox{DC}}(m,n;\;\hat{C}_{8})}{h_{8,\scriptsize\mbox{DC}}(\cdot,\cdot;\;\hat{C}_{8})*1|_{(x,y)^{T}=(0,0)^{T}}},\end{split}$ (73)
so that we obtain the following desirable properties
$\begin{split}h_{7,\scriptsize\mbox{norm}}(m,n)\star 1=1,\end{split}$ (74)
$\begin{split}h_{8,\scriptsize\mbox{norm}}(m,n)\star 1=1,\end{split}$ (75)
in analogy to corresponding relationships for continuous Gaussian kernels and appropriate discrete approximations thereof.
$\sqrt{\det V(|h_{7}(\cdot,\cdot)-C_{7}|)}$
$\sqrt{\det V(|h_{8}(\cdot,\cdot)-C_{8}|)}$
Figure 5: Graphs of the error measure $\sqrt{\det V(|h_{i}(\cdot,\cdot)-C_{i}|)}$ for $i\in\{7,8\}$ when determining the DC compensation constants $C_{7}\approx-0.0118$ and $C_{8}\approx-0.0386$ according to (71), in order to later renormalize the Gaussian-like “master key filters” to unit $l_{1}$-norm, as regular Gaussian kernels obey.
Filter $\|h_{i}(\cdot,\cdot)\|_{1}$ $M(|h_{i,\scriptsize\mbox{DC}}(\cdot,\cdot))|$ $V(|h_{i,\scriptsize\mbox{DC}}(\cdot,\cdot))|$
7 2.686 $\left(\begin{array}[]{r}+0.050\\ -0.021\end{array}\right)$ $\left(\begin{array}[]{rr}+0.600&-0.018\\ -0.018&+0.962\end{array}\right)$
8 1.055 $\left(\begin{array}[]{r}+0.019\\ +0.028\end{array}\right)$ $\left(\begin{array}[]{rr}+0.696&+0.028\\ -0.018&+0.680\end{array}\right)$
Table 4: The result of computing the discrete $l_{1}$-norm $\|h(\cdot,\cdot)\|_{1}$ as well as the discrete spatial mean value $M(|h(\cdot,\cdot))|$ according to Equation (3) and the discrete spatial variance $V(|h(\cdot,\cdot))|$ according to Equation (4) for the DC-compensated “master key filters” 7 and 8 according to (69) and (70).
Filter $h_{i,\scriptsize\mbox{norm}}(\cdot,\cdot)\star 1$ $h_{i,\scriptsize\mbox{norm}}(\cdot,\cdot)\star(x-m_{x})$ $h_{i,\scriptsize\mbox{norm}}(\cdot,\cdot)\star(y-m_{y})$
7 1.000 -0.098 -0.272
8 1.000 -0.057 -0.117
Table 5: The result of computing the responses to the monomials $\theta_{00}=1$, $\theta_{10}=x-m_{x}$ and $\theta_{01}=y-m_{y}$ for normalized versions of the last 2 “master key filters” shown in Figure 1, where the spatial offset $(m_{x},m_{y})^{T}=M(|h(\cdot,\cdot)|$ represents the spatial mean of the absolute value of each filter. For this table, the amplitudes of the original last 2 “master key filters” have been rescaled, to make the response equal to 1 for the monomial of $\theta_{00}=1$ equal to 1.
Figure 5 show graphs of the error measure
$\sqrt{\det V(|h_{i}(\cdot,\cdot)-C_{i}|)}$ (76)
according to (71) for determining the DC compensation constants for $\hat{C_{7}}$ and $\hat{C_{8}}$ in this way. Table 4 shows the result of computing the spatial spread measures for the resulting DC-compensated filters $h_{7,\scriptsize\mbox{DC}}(m,n)$ and $h_{8,\scriptsize\mbox{DC}}(m,n)$ according to (69) and (70). As can be seen from the results, the use of DC compensation substantially reduces the variance-based spatial spread measures for Filters 7 and 8, compared to the variance-based spread measures of the corresponding not DC-compensated filters shown in Table 1.
Table 5 additionally shows the responses to monomials up to order 1 for the resulting DC-compensated and $l_{1}$-normalized filters $h_{7,\scriptsize\mbox{norm}}(m,n)$ and $h_{8,\scriptsize\mbox{norm}}(m,n)$ according to (72) and (73). As can be seen from the table, for the larger-size Gaussian-like Filter 8, the responses to the first-order monomials $x$ and $y$ are rather low. For the smaller-size Gaussian-like Filter 7, the deviations of those responses from the ideal value 0 are, however, notable.
Figure 2 shows a visualization of the resulting normalized versions of the 8 “master key filters”, which we will then approximate by idealized receptive field models in different ways in Section 3. Before proceeding to the modelling stage, we will, however, first in the next section introduce a way of defining spatial spread measures, that are less influenced by spurious variabilities in the peripheries of the learned filters.
2.5 Reducing the bias in the estimates of the spatial variances, by windowed estimates of the spatial spread measures
Due to spurious variations in the background of the learned filters, these spurious variations may lead to substantial biases in the mean-based and variance-based spread measures, because those spurious variations are weighted with either a linear or a quadratic factor in the first- and second-order moments that define the spatial spread measures. This situation is, thus, in contrast to previous use of such spatial spread measures in Lindeberg (2024, 2025) for characterizing properties of idealized mathematical models of receptive fields with regard to variabilities in their spatial extent under variations of the spatial scale parameter.
2.5.1 Weighted spatial spread measures
To reduce this substantial source to bias, we will in this section replace the previous uniformly weighted first- and second-order moments by weighted first- and second-order moments, with the weight functions determined from discrete approximations of Gaussian functions with similar values of the scale parameters as we would obtain by estimating the scale parameters of idealized filter models from the spatial variances of the corresponding discrete kernels. Thus, we define weighted spatial mean values and spatial variance of continuous or discrete filter kernels $h(\cdot,\cdot)$ according to:
$M_{g(\cdot)}(h(\cdot))=\frac{\int_{p\in{\mathbb{R}}^{2}}p\,h(p)\,g(p)\,dp}{\int_{p\in{\mathbb{R}}^{2}}h(p)\,g(p)\,dp},$ (77)
$V_{g(\cdot)}(h(\cdot))=\frac{\int_{p\in{\mathbb{R}}^{2}}p\,p^{T}\,h(p)\,g(p)\,dp}{\int_{p\in{\mathbb{R}}^{2}}h(p)\,g(p)\,dp}\\ -\left(\frac{\int_{p\in{\mathbb{R}}^{2}}p\,h(p)\,g(p)\,dp}{\int_{p\in{\mathbb{R}}^{2}}h(p)\,g(p)\,dp}\right)^{2},$ (78)
over a continuous image domain and according to
$M_{T(\cdot)}(h(\cdot))=\frac{\sum_{p\in{\mathbb{R}}^{2}}p\,h(p)\,T(p)\,dp}{\sum_{p\in{\mathbb{R}}^{2}}h(p)\,T(p)\,dp},$ (79)
$V_{T(\cdot)}((h(\cdot))=\frac{\sum_{p\in{\mathbb{R}}^{2}}p\,p^{T}\,h(p)\,T(p)}{\sum_{p\in{\mathbb{R}}^{2}}h(p)\,T(p)}\\ -\left(\frac{\sum_{p\in{\mathbb{R}}^{2}}p\,h(p)\,T(p)}{\sum_{p\in{\mathbb{R}}^{2}}h(p)\,T(p)}\right)^{2},$ (80)
over a discrete image domain, with the intention that peripheral values corresponding to spurious structures in the learned filters should have lower influence on the spatial spread measures.
Specifically, if the continuous receptive field $h(\cdot)$ would be equal to the continuous weighting filter $g(\cdot)$, then the weighted spatial mean value and the weighted spatial variance would correspond to mean value and standard deviation in $L_{2}$-norm. Similarly, if the discrete receptive field $h(\cdot)$ would be equal to the discrete weighting filter $T(\cdot)$, then the weighted spatial mean value and the weighted spatial variance would correspond to mean value and standard deviation in $l_{2}$-norm. For our purpose, we do other hand formulate a hybrid approach, where we use the desired shape of the filter model as a prior to suppress the influence of spurious responses in the learned filter spatial positions in the peripheries of the support regions of the learned filters.
2.5.2 Determining idealized shapes for the weighting kernels from the qualitative shapes of the “master key filters”
The intention is then to choose either continuous or discrete weighting filters that match the qualitative shapes of the filters to be modelled in the following ways:
  • For Filter 8 we will either use anisotropic continuous Cartesian-aligned anisotropic Gaussian kernels
    $g(x,y;\;\sigma_{x},\sigma_{y})=g_{\scriptsize\mbox{1D}}(x;\;\sigma_{x})\,g_{\scriptsize\mbox{1D}}(y;\;\sigma_{y})$ (81)
    or discrete Cartesian-aligned anisotropic Gaussian kernels
    $T(m,n;\;\sigma_{x},\sigma_{y})=T_{\scriptsize\mbox{1D}}(m;\;\sigma_{x})\,T_{\scriptsize\mbox{1D}}(n;\;\sigma_{y})$ (82)
    having possibly different values of the scale parameters $\sigma_{x}$ and $\sigma_{y}$ in the $x$- and $y$-directions, respectively.
  • For Filters 2, 4 and 5, we will either use the absolute value of either a first-order horizontal $x$-derivative of a continuous Gaussian kernel
    $g_{x}(\cdot,\cdot;\;\sigma_{x},\sigma_{y})=g_{\scriptsize\mbox{1D},x}(x;\;\sigma_{x})\,g_{\scriptsize\mbox{1D}}(y;\;\sigma_{y})$ (83)
    or the absolute value of a discrete approximation of a horizontal $x$-derivative of a Gaussian kernel
    $(\Delta_{x}T)(\cdot,\cdot;\;\sigma_{x},\sigma_{y})=(\Delta_{x}T_{\scriptsize\mbox{1D}})(x;\;\sigma_{x})\,T_{\scriptsize\mbox{1D}}(y;\;\sigma_{y})$ (84)
    for appropriately matched difference operators $\Delta_{x}\in\{\delta_{x-},\delta_{x+},\delta_{x}\}$ in the horizontal $x$-direction, with the scale parameters $\sigma_{x}$ and $\sigma_{y}$ determined from variance-based spatial spread measures of the corresponding filters.
  • For Filters 1, 3 and 6, we will either use the absolute value of a first-order vertical $y$-derivative of continuous Gaussian kernel
    $g_{y}(\cdot,\cdot;\;\sigma_{x},\sigma_{y})=g_{\scriptsize\mbox{1D}}(x;\;\sigma_{x})\,g_{\scriptsize\mbox{1D},y}(y;\;\sigma_{y})$ (85)
    or the absolute value of a discrete approximation of a vertical $y$-derivative of a Gaussian kernel
    $(\Delta_{y}T)(\cdot,\cdot;\;\sigma_{x},\sigma_{y})=T_{\scriptsize\mbox{1D}}(x;\;\sigma_{x})\,(\Delta_{y}T_{\scriptsize\mbox{1D}})(y;\;\sigma_{y})$ (86)
    for a appropriately matched difference operators $\Delta_{y}\in\{\delta_{y+},\delta_{y-},\delta_{y}\}$ in the vertical $y$-direction, with the scale parameters $\sigma_{x}$ and $\sigma_{y}$ determined from variance-based spatial spread measures of the corresponding filters.
The motivation for these choices is to use qualitatively similar shapes for the spatial weighting functions in the spatial spread measures as the absolute values of the corresponding idealized models of receptive field shapes.
2.6 Estimating variance parameters of receptive fields from weighted spatial spread measures
When using non-uniform weighting kernels when computing the spatial spread measures, the resulting variance-based measures will, however, be affected by the spatial extent of the weighting kernels. In the continuous case, we specifically have the following relations:
$\begin{split}M_{g(\cdot,\cdot;\;\sigma_{x},\sigma_{y})}(g(\cdot,\cdot;\;\sigma_{x},\sigma_{y}))=\left(\begin{array}[]{c}0\\ 0\end{array}\right),\end{split}$ (87)
$\begin{split}V_{g(\cdot,\cdot;\;\sigma_{x},\sigma_{y})}(g(\cdot,\cdot;\;\sigma_{x},\sigma_{y}))=\left(\begin{array}[]{cc}\sigma_{x}^{2}/2&0\\ 0&\sigma_{y}^{2}/2\end{array}\right),\end{split}$ (88)
$\begin{split}M_{|g_{x}(\cdot,\cdot;\;\sigma_{x},\sigma_{y})|}(|g_{x}(\cdot,\cdot;\;\sigma_{x},\sigma_{y})|)=\left(\begin{array}[]{c}0\\ 0\end{array}\right),\end{split}$ (89)
$\begin{split}V_{|g_{x}(\cdot,\cdot;\;\sigma_{x},\sigma_{y})|}(|g_{x}(\cdot,\cdot;\;\sigma_{x},\sigma_{y})|)=\left(\begin{array}[]{cc}3\,\sigma_{x}^{2}/2&0\\ 0&\sigma_{y}^{2}/2\end{array}\right),\end{split}$ (90)
$\begin{split}M_{|g_{y}(\cdot,\cdot;\;\sigma_{x},\sigma_{y})|}(|g_{y}(\cdot,\cdot;\;\sigma_{x},\sigma_{y})|)=\left(\begin{array}[]{c}0\\ 0\end{array}\right),\end{split}$ (91)
$\begin{split}V_{|g_{y}(\cdot,\cdot;\;\sigma_{x},\sigma_{y})|}(|g_{y}(\cdot,\cdot;\;\sigma_{x},\sigma_{y})|)=\left(\begin{array}[]{cc}\sigma_{x}^{2}/2&0\\ 0&3\,\sigma_{y}^{2}/2\end{array}\right).\end{split}$ (92)
Our intention is then to use these expressions for the weighted variance-based spatial spread measures for estimating the scale parameters $\sigma_{x}$ and $\sigma_{y}$ in idealized models of the learned “master key filters”, based on weighted variance measurements for the learned filters.
Filter $M_{T(\cdot,\cdot)}(|h_{i}(\cdot,\cdot))|$ $V_{T(\cdot,\cdot)}(|h_{i}(\cdot,\cdot))|$
1 $\left(\begin{array}[]{r}-0.005\\ -0.753\end{array}\right)$ $\left(\begin{array}[]{rr}+0.169&-0.010\\ -0.010&+0.466\end{array}\right)$
2 $\left(\begin{array}[]{r}+0.665\\ -0.017\end{array}\right)$ $\left(\begin{array}[]{rr}+0.469&+0.010\\ +0.010&+0.168\end{array}\right)$
3 $\left(\begin{array}[]{r}+0.011\\ +0.666\end{array}\right)$ $\left(\begin{array}[]{rr}+0.190&+0.001\\ +0.001&+0.541\end{array}\right)$
4 $\left(\begin{array}[]{r}-0.536\\ +0.005\end{array}\right)$ $\left(\begin{array}[]{rr}+0.619&+0.000\\ +0.000&+0.090\end{array}\right)$
5 $\left(\begin{array}[]{r}-0.011\\ -0.020\end{array}\right)$ $\left(\begin{array}[]{rr}+1.732&-0.011\\ -0.011&+0.320\end{array}\right)$
6 $\left(\begin{array}[]{r}-0.006\\ +0.114\end{array}\right)$ $\left(\begin{array}[]{rr}+0.297&-0.002\\ -0.002&+1.510\end{array}\right)$
8 $\left(\begin{array}[]{r}-0.002\\ +0.008\end{array}\right)$ $\left(\begin{array}[]{rr}+0.152&+0.001\\ +0.001&+0.149\end{array}\right)$
Table 6: The result of computing the discrete weighted spatial mean value $M_{T(\cdot,\cdot)}(|h(\cdot,\cdot))|$ according to Equation (79) and the discrete weighted spatial variance $V_{T(\cdot,\cdot)}(|h(\cdot,\cdot))|$ according to Equation (80) for the normalized versions of the 8 “master key filters” according to (63)–(68) and (73) when using matching shapes of the spatial weighting filters $T(\cdot,\cdot)$ to the qualitative shapes of the “master key filters” according to Section 2.5.2.
Table 6 shows weighted spatial spread measures computed in this way, with the weighting functions for the respective normalized versions of the “master key filters” according to (63)–(68) and (73) determined from the absolute values of the corresponding idealized filter models according to (93)–(100), with the scale parameters of the weighting functions set to $\sigma_{x}=\sigma_{y}=\sigma_{0}=1$ in the spatial weighting kernels for all the “master key filters”.
As can be seen from a comparison with the corresponding non-weighted spatial spread measures in Tables 1 and 4, the use of spatial weighting in the computation of the variance-based spread measures lead to substantially lower estimates of the corresponding scale parameters, and thus to a substantial decrease of the previous bias, caused by spurious variations in the peripheries of the learned filters.
We can also see that the elements of the weighted mean value vectors are, as for the previously computed non-weighted mean values, either around $\pm 1/2$ or 0, although again with slight offsets from the ideal values by up to 0.1 or 0.2 grid units. In this respect, both the weighted and the unweighted mean value give rise to qualitatively similar results, indicating that the spurious variations in the backgrounds of the “master key filters” seem to have roughly symmetric statistical properties under mirroring through the coordinate axes.
3 Modelling the 8 “master key filters” in terms of idealized scale-space operations
In the following, we will use the modelling and analysis results from the previous section, for modelling the 8 “master key filters” in terms of idealized scale-space operations. Specifically, we will explore the implications of using different types of modelling criteria and analyze how they give rise to different sets of scale parameters in the idealized filter models. Overall, we will either consider separate modelling of the different filters of the following forms:
$\begin{split}h_{1,\scriptsize\mbox{ideal}}(m,n;\;\sigma_{x,1},\sigma_{y,1})=(\delta_{y+}T)(m,n;\;\sigma_{x,1},\sigma_{y,1}),\end{split}$ (93)
$\begin{split}h_{2,\scriptsize\mbox{ideal}}(m,n;\;\sigma_{x,2},\sigma_{y,2})=(\delta_{x-}T)(m,n;\;\sigma_{x,2},\sigma_{y,2}),\end{split}$ (94)
$\begin{split}h_{3,\scriptsize\mbox{ideal}}(m,n;\;\sigma_{x,3},\sigma_{y,3})=(\delta_{y-}T)(m,n;\;\sigma_{x,3},\sigma_{y,3}),\end{split}$ (95)
$\begin{split}h_{4,\scriptsize\mbox{ideal}}(m,n;\;\sigma_{x,4},\sigma_{y,4})=(\delta_{x+}T)(m,n;\;\sigma_{x,4},\sigma_{y,4}),\end{split}$ (96)
$\begin{split}h_{5,\scriptsize\mbox{ideal}}(m,n;\;\sigma_{x,5},\sigma_{y,5})=(\delta_{x}T)(m,n;\;\sigma_{x,5},\sigma_{y,5}),\end{split}$ (97)
$\begin{split}h_{6,\scriptsize\mbox{ideal}}(m,n;\;\sigma_{x,6},\sigma_{y,6})=(\delta_{y}T)(m,n;\;\sigma_{x,6},\sigma_{y,6}),\end{split}$ (98)
$\begin{split}h_{7,\scriptsize\mbox{ideal}}(m,n;\;\sigma_{x,7},\sigma_{y,7})=1-\gamma_{7}\,(\nabla_{5}^{2}T)(m,n;\;\sigma_{7},\sigma_{7}),\end{split}$ (99)
$\begin{split}h_{8,\scriptsize\mbox{ideal}}(m,n;\;\sigma_{x,8},\sigma_{y,8})=T(m,n;\;\sigma_{x,8},\sigma_{y,8}),\end{split}$ (100)
with $T(m,n;\;\sigma_{x},\sigma_{y})$ denoting a 2-D discrete Gaussian kernel with possibly different scale parameters $\sigma_{x}$ and $\sigma_{y}$ in the two coordinate directions according to (82), and for the special case of Filter 7, the operator $\nabla_{5}^{2}=\delta_{xx}+\delta_{yy}$ denoting a discrete approximation of the Laplacian operator, and $\gamma_{7}\in{\mathbb{R}}_{+}$ denoting a constant to be determined. To keep the dimensionality of the search space down, as well as for theoretical reasons, we have here stated an a priori spatially isotropic model for Filter 7 with $\sigma_{x,7}=\sigma_{y,7}=\sigma_{7}$.
The overall form of modelling above is based on the canonical model for spatial receptive fields of the form (Lindeberg 2021a Equation (23))
$T_{{\varphi}^{m_{1}}{\bot{\varphi}}^{m_{2}}}(x,y;\;s,\Sigma)=\partial_{\varphi}^{m_{1}}\partial_{\bot\varphi}^{m_{2}}\left(g(x,y;\;s\Sigma)\right),$ (101)
while here (i) restricting the spatial covariance matrix $\Sigma$ in the affine Gaussian kernel $g(x,y;\;s\Sigma)$ to a diagonal matrix according to $\Sigma_{i}=\operatorname{diag}(\sigma_{x,i}^{2},\sigma_{y,i}^{2})$ for each filter $i$ and (ii) restricting the directional derivative operators $\partial_{\varphi}$ and $\partial_{\bot\varphi}$ to Cartesian partial derivatives $\partial_{x}$ and $\partial_{y}$, since the spatial shapes of the learned receptive fields appear to be very much aligned with the Cartesian coordinate directions.
Then, we also make use of the fact that the discrete analogue of the Gaussian kernel in (82) and (9) constitutes a canonical way to transfer the scale-space properties of the continuous Gaussian kernel to a discrete domain. The use of complementary centered or non-centered difference operators in (93)–(100) is based on the property that such difference operators applied to a discrete scale-space representation generated by convolution with the discrete analogue of the Gaussian kernel preserve discrete scale-space properties (Lindeberg 1993a,2024).
One could, however, also conceive using other discrete approximations of Gaussian derivative operators, such as those in (Lindeberg 2024, 2025)
In the following Sections 3.13.6 we will consider different ways of determining the scale parameters $\sigma_{x,i}$ and $\sigma_{y,i}$ for $i\in\{1,2,3,4,5,6,8\}$, while we will then in Section 3.7 consider a different methodology with one more parameter for determining the parameters $\gamma_{7}$ and $\sigma_{7}$ in an idealized model of Filter 7.
Note:
In the visualizations for the resulting Methods A–B to be presented, we will, however, not develop specific variance-based methods for Filter 7, since the idealized model of that filter also comprises an amplitude scaling factor $\gamma_{7}$ for the contribution from the negative Laplacian-of-the-Gaussian. For this reason, we will for Methods A–B instead interweave visualizations of idealized models of Filter 7 based on the either $l_{2}$-norm-based or $l_{1}$-norm-based Methods D2 and C2, to be described in Section 3.7.
Filter $\hat{\sigma}_{x,i}$ $\hat{\sigma}_{y,i}$
1 0.580 0.558
2 0.559 0.580
3 0.617 0.601
4 0.642 0.424
5 1.075 0.800
6 0.771 1.003
8 0.552 0.545
Table 7: Estimates of the horizontal and vertical scale parameters $\hat{\sigma}_{x,i}$ and $\hat{\sigma}_{y,i}$ computed according to Method A, based on estimates from continuous Gaussian derivatives according to (103)–(116) in idealized models of the “master key filters” according to (93)–(100), in turn based on weighted measurements of the variance of the absolute values of the learned filters according to Table 6.
3.1 Method A: Direct transfer of scale values from spatial variances to idealized scale-space operations with different amounts of smoothing in the different coordinate directions with use of continuous Gaussian derivative model
A straightforward way of estimating the scale parameters $\sigma_{x,i}$ and $\sigma_{y,i}$ in the idealized models (93)–(100) is by making use of the closed-form expressions for the weighted variances of the corresponding continuous Gaussian derivative kernels according to (88), (90) and (92). Denoting the weighted variance matrix for each continuous kernel by
$V_{|g_{\alpha}(\cdot,\cdot;\;\sigma_{x},\sigma_{y})|}(|g_{\alpha}(\cdot,\cdot;\;\sigma_{x},\sigma_{y})|)=\left(\begin{array}[]{cc}v_{xx}&v_{xy}\\ v_{xy}&v_{yy}\end{array}\right),$ (102)
we then obtain the following estimates
$\begin{split}\hat{\sigma}_{x,1}=\sqrt{2\,v_{xx,1}},\end{split}$ (103)
$\begin{split}\hat{\sigma}_{y,1}=\sqrt{2\,v_{yy,1}/3},\end{split}$ (104)
$\begin{split}\hat{\sigma}_{x,2}=\sqrt{2\,v_{xx,2}/3},\end{split}$ (105)
$\begin{split}\hat{\sigma}_{y,2}=\sqrt{2\,v_{yy,2}},\end{split}$ (106)
$\begin{split}\hat{\sigma}_{x,3}=\sqrt{2\,v_{xx,3}},\end{split}$ (107)
$\begin{split}\hat{\sigma}_{y,3}=\sqrt{2\,v_{yy,3}/3},\end{split}$ (108)
$\begin{split}\hat{\sigma}_{x,4}=\sqrt{2\,v_{xx,4}/3},\end{split}$ (109)
$\begin{split}\hat{\sigma}_{y,4}=\sqrt{2\,v_{yy,4}},\end{split}$ (110)
$\begin{split}\hat{\sigma}_{x,5}=\sqrt{2\,v_{xx,5}/3},\end{split}$ (111)
$\begin{split}\hat{\sigma}_{y,5}=\sqrt{2\,v_{yy,5}},\end{split}$ (112)
$\begin{split}\hat{\sigma}_{x,6}=\sqrt{2\,v_{xx,6}},\end{split}$ (113)
$\begin{split}\hat{\sigma}_{y,6}=\sqrt{2\,v_{yy,6}/3},\end{split}$ (114)
$\begin{split}\hat{\sigma}_{x,8}=\sqrt{2\,v_{xx,8}},\end{split}$ (115)
$\begin{split}\hat{\sigma}_{y,8}=\sqrt{2\,v_{yy,8}},\end{split}$ (116)
with numerical values of these entities shown in Table 7 based on the numerical values of the variances in Table 6.
Figure 3 shows visualizations of the corresponding idealized receptive fields.
$\begin{split}\hat{\sigma}_{x,i}=\operatorname{argmin}_{\sigma_{x,i}}|V_{xx,|h_{i,\scriptsize\mbox{ideal}}(\cdot,\cdot;\;\sigma_{0},\sigma_{0})|}(|h_{i,\scriptsize\mbox{ideal}}(\cdot,\cdot;\;\sigma_{x,1},\sigma_{y,1})|)-V_{xx,|h_{i,\scriptsize\mbox{ideal}}(\cdot,\cdot;\;\sigma_{0},\sigma_{0})|}(|h_{i,\scriptsize\mbox{norm}}(\cdot,\cdot)|)|\end{split}$ (117)
$\begin{split}\hat{\sigma}_{y,i}=\operatorname{argmin}_{\sigma_{y,i}}|V_{yy,|h_{i,\scriptsize\mbox{ideal}}(\cdot,\cdot;\;\sigma_{0},\sigma_{0})|}(|h_{i,\scriptsize\mbox{ideal}}(\cdot,\cdot;\;\sigma_{x,1},\sigma_{y,1})|)-V_{yy,|h_{i,\scriptsize\mbox{ideal}}(\cdot,\cdot;\;\sigma_{0},\sigma_{0})|}(|h_{i,\scriptsize\mbox{norm}}(\cdot,\cdot)|)|\end{split}$ (118)
Figure 6: Formal specification of the criteria used for determining the scale estimates $\hat{\sigma}_{x,i}$ and $\hat{\sigma}_{y,i}$ in the horizontal and vertical directions, by matching the diagonal elements of the discrete variance-based spatial spread measures between an idealized receptive field model $h_{i,\scriptsize\mbox{ideal}}(\cdot,\cdot;\;\sigma_{x,1},\sigma_{y,1})$ according to (93)–(131) and (100) to the corresponding discrete variance-based spatial spread measures for the normalized versions $h_{i,\scriptsize\mbox{norm}}(\cdot,\cdot)|)$ of the “master key filters” according to (63)–(68) and (73). The parameter $\sigma_{0}$ specifies the spatial extent if the discrete weighting kernel used for computing the weighted variance-based spatial spread measures, and is provided as prior information, chosen as $\sigma_{0}=1$ for all the filters. (Note that since the idealized receptive field models $h_{i,\scriptsize\mbox{ideal}}(\cdot,\cdot;\;\sigma_{x,1},\sigma_{y,1})$ are separable with respect to the Cartesian coordinate directions $x$ and $y$, the vertical parameter $\sigma_{y,1}$ does influence the estimate of the horizontal diagonal element $V_{xx,|h_{i,\scriptsize\mbox{ideal}}(\cdot,\cdot;\;\sigma_{0},\sigma_{0})|}(|h_{i,\scriptsize\mbox{ideal}}(\cdot,\cdot;\;\sigma_{x,1},\sigma_{y,1})|)$. Similarly, the horizontal parameter $\sigma_{x,1}$ does not influence the estimate of the vertical diagonal element $V_{yy,|h_{i,\scriptsize\mbox{ideal}}(\cdot,\cdot;\;\sigma_{0},\sigma_{0})|}(|h_{i,\scriptsize\mbox{ideal}}(\cdot,\cdot;\;\sigma_{x,1},\sigma_{y,1})|)$.)
Idealized 1(B) Idealized 2(B) Idealized 3(B) Idealized 4(B)
Refer to caption Refer to caption Refer to caption Refer to caption
Idealized 5(B) Idealized 6(B) Idealized 7(C2) Idealized 8(B)
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 7: Idealized models of the 8 “master key filters” in Figure 2, in terms of combinations pure difference operators with spatial smoothing filters based on the discrete analogue of the Gaussian kernel, with the amount of spatial smoothing determined from the spatial extents of the 8 “master key filters” using Method B in Section 3.2, based on requiring the horizontal vs. the vertical discrete weighted variance-based spatial spread measures for the idealized receptive field models to be equal to the weighted discrete spatial spread measures for the corresponding learned filters. (Idealized 1): Non-centered vertical difference operator $(\delta_{y+}T)(m,n;\;\sigma_{x,1},\sigma_{y,1})$. (Idealized 2): Non-centered horizontal difference operator $(\delta_{x-}T)(m,n;\;\sigma_{x,2},\sigma_{y,2})$. (Idealized 3): Non-centered vertical difference operator $(\delta_{y-}T)(m,n;\;\sigma_{x,3},\sigma_{y,3})$. (Idealized 4): Non-centered horizontal difference operator $(\delta_{x+}T)(m,n;\;\sigma_{x,4},\sigma_{y,4})$. (Idealized 5): Centered horizontal difference operator $(\delta_{x}T)(m,n;\;\sigma_{x,5},\sigma_{y,5})$. (Idealized 6): Centered vertical difference operator $(\delta_{y}T)(m,n;\;\sigma_{x,6},\sigma_{y,6})$. (Idealized 7): Sharpening operation $1-\gamma_{7}\,(\nabla_{5}^{2}T)(m,n;\;\sigma_{7},\sigma_{7})$. (Idealized 8): Pure spatial smoothing $T(m,n;\;\sigma_{x,8},\sigma_{y,8})$. (The numerical values of the scale parameters $\sigma_{x,i}$ and $\sigma_{y,i}$ for $i\in\{1,2,3,4,5,6,8\}$ have been determined according to Table 8 and with $\sigma_{7}$ additionally according to Table 13.) (Note that the contrasts of Filters 2 and 3 are reversed in relation to the sign conventions for corresponding “master key filters”.)
3.2 Method B: Matching of discrete spatial variances between the “master key filters” and the spatial variances of discrete derivative approximation filters, with individual scale parameters for each filter
A possible limitation of the above computed estimates of the spatial scale parameters from weighted variances of the discrete filters, is that the relationship between the weighted variances of the filters is based on a continuous model in terms of Gaussian derivatives (Equations (88), (90) and (92) for continuous Gaussian derivatives, that lead to the estimates in Equations (103)–(116)), while the filters that we model are genuinely discrete, and specifically corresponding to rather fine scale levels, where discretization effects in relation to the discrete vs. continuous models may therefore be significant.
To address that issue, we will in this section instead consider a fully discrete approach, where discrete weighted variances of fully discrete models of the idealized receptive fields are matched to discretely computed weighted variances of the learned filters.
Denoting the operators that extract the diagonal elements in the discrete weighted variance-based spatial spread measure $V_{T(\cdot)}(h(\cdot))$ in (4) by $V_{xx,T(\cdot)}(h(\cdot))$ and $V_{yy,T(\cdot)}(h(\cdot))$, respectively, we will in the following determine the scale estimates in idealized models of the “master key filters” according to (93)–(100), based on the criterion that the differences between the diagonal elements of the variance-based spatial spread measures should be as small as possible, between the idealized receptive field models and the learned filters, as specified in Equations (117)–(118) in Figure 6.
Filter $\hat{\sigma}_{x,i}$ $\hat{\sigma}_{y,i}$
1 0.644 0.583
2 0.586 0.644
3 0.690 0.674
4 0.756 0.460
5 1.107 0.945
6 0.900 0.889
8 0.609 0.601
Table 8: Estimates of the horizontal and vertical scale parameters $\hat{\sigma}_{x,i}$ and $\hat{\sigma}_{y,i}$ computed according to Method B, based on requiring the horizontal vs. the vertical discrete weighted variance-based spatial spread measures for the idealized models of the “master key filters” to be equal to the weighted discrete spatial spread measures for the the corresponding learned filters.
Table 8 shows the result of computing horizontal and vertical scale estimates $\hat{\sigma}_{x,i}$ and $\hat{\sigma}_{x,i}$ in this way for the 8 different “master key filters”.
Figure 7 shows visualizations of the corresponding idealized receptive field models of these filters.
Idealized 1(C1) Idealized 2(C1) Idealized 3(C1) Idealized 4(C1)
Refer to caption Refer to caption Refer to caption Refer to caption
Idealized 5(C1) Idealized 6(C1) Idealized 7(C2) Idealized 8(C1)
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 8: Idealized models of the 8 “master key filters” in Figure 2, in terms of combinations pure difference operators with spatial smoothing filters based on the discrete analogue of the Gaussian kernel, with the amount of spatial smoothing determined from the spatial extents of the 8 “master key filters” using Method C1 in Section 3.3, based on minimizing the discrete $l_{1}$-norm between the idealized receptive field models and the normalized versions of the learned filters and using different values of the scale parameters in the horizontal and the vertical directions according to (121). (Idealized 1): Non-centered vertical difference operator $(\delta_{y+}T)(m,n;\;\sigma_{x,1},\sigma_{x,1})$. (Idealized 2): Non-centered horizontal difference operator $(\delta_{x-}T)(m,n;\;\sigma_{x,2},\sigma_{x,2})$. (Idealized 3): Non-centered vertical difference operator $(\delta_{y-}T)(m,n;\;\sigma_{x,3},\sigma_{x,3})$. (Idealized 4): Non-centered horizontal difference operator $(\delta_{x+}T)(m,n;\;\sigma_{x,4},\sigma_{x,4})$. (Idealized 5): Centered horizontal difference operator $(\delta_{x}T)(m,n;\;\sigma_{x,5},\sigma_{x,5})$. (Idealized 6): Centered vertical difference operator $(\delta_{y}T)(m,n;\;\sigma_{x,6},\sigma_{x,6})$. (Idealized 7): Sharpening operation $1-\gamma_{7}\,(\nabla_{5}^{2}T)(m,n;\;\sigma_{7},\sigma_{7})$. (Idealized 8): Pure spatial smoothing $T(m,n;\;\sigma_{x,8},\sigma_{x,8})$. (The numerical values of $\sigma_{i}$ for $i\in\{1,2,3,4,5,6,8\}$ are according to Table 9, and with $\sigma_{7}$ additionally according to Table 13.) (Note that the contrasts of Filters 2 and 3 are reversed in relation to the sign conventions for corresponding “master key filters”.)
Filter $\hat{\sigma}_{x,i}$ $\hat{\sigma}_{y,i}$
1 0.360 0.510
2 0.555 0.453
3 0.701 0.655
4 0.563 0.384
5 1.309 0.875
6 0.973 1.171
8 0.637 0.587
Table 9: Estimates of the horizontal and vertical scale parameters $\hat{\sigma}_{x,i}$ and $\hat{\sigma}_{y,i}$ computed according to Method C1, based on minimizing the discrete $l_{1}$-norm between the idealized receptive field model $h_{i,\scriptsize\mbox{ideal}}(\cdot,\cdot;\;\sigma_{x,1},\sigma_{y,1})$ according to (93)–(100) and the normalized learned filter $h_{i,\scriptsize\mbox{norm}}(\cdot,\cdot)$ according to (63)–(68) and (73).
3.3 Method C1: $l_{1}$-norm-based fitting to idealized model for scale-space operations with different scale parameters along the horizontal and the vertical directions
Although the above spatial spread-measure based modelling operations aim at reflecting the spatial extents of the receptive fields, a conceptual problem arises if there is a mismatch between the shapes of the filters in the idealized receptive fields to the shapes of the learned receptive fields.
In this section, we will address this issue by instead fitting the idealized receptive field models to the learned receptive fields in discrete $l_{1}$-norm. Thus, we will estimate the scale parameters of the receptive fields in the following way:
$(\hat{\sigma}_{x,i},\hat{\sigma}_{y,i})=\operatorname{argmin}_{(\hat{\sigma}_{x,i},\hat{\sigma}_{y,i})}\|\Delta h_{i,\scriptsize\mbox{norm}}(\cdot,\cdot;\;\sigma_{x,i},\sigma_{y,i})\|_{1}=\\ =\operatorname{argmin}_{(\hat{\sigma}_{x,i},\hat{\sigma}_{y,i})}\|h_{i,\scriptsize\mbox{ideal}}(\cdot,\cdot;\;\sigma_{x,i},\sigma_{y,i})-h_{i,\scriptsize\mbox{norm}}(\cdot,\cdot)\|_{1}$ (119)
with the idealized receptive field models $h_{i,\scriptsize\mbox{ideal}}(\cdot,\cdot;\;\sigma_{x,1},\sigma_{y,1})$ according to (93)–(100) and the normalized “master key filters” $h_{i,\scriptsize\mbox{norm}}(\cdot,\cdot)$ according to (63)–(68) and (73).
Table 9 shows the resulting horizontal and vertical scale estimates obtained in this way. Figure 8 shows visualizations of the corresponding idealized receptive fields.
Filter $\hat{\sigma}_{i}$
1 0.458
2 0.448
3 0.671
4 0.420
5 1.387
6 1.090
8 0.611
Table 10: Estimates of the joint horizontal and vertical scale parameters $\hat{\sigma}_{i}$ computed according to Method C2, based on minimizing the discrete $l_{1}$-norm between the idealized receptive field model $h_{i,\scriptsize\mbox{ideal}}(\cdot,\cdot;\;\sigma_{x,1},\sigma_{y,1})$ according to (93)–(100) and the normalized learned filter $h_{i,\scriptsize\mbox{norm}}(\cdot,\cdot)$ according to (63)–(68) and (73), under the assumption that the values of the horizontal and the vertical scale parameters are the same $\sigma_{x,i}=\sigma_{y,i}=\sigma_{i}$.
Idealized 1(C2) Idealized 2(C2) Idealized 3(C2) Idealized 4(C2)
Refer to caption Refer to caption Refer to caption Refer to caption
Idealized 5(C2) Idealized 6(C2) Idealized 7(C2) Idealized 8(C2)
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 9: Idealized models of the 8 “master key filters” in Figure 2, in terms of combinations pure difference operators with spatial smoothing filters based on the discrete analogue of the Gaussian kernel, with the amount of spatial smoothing determined from the spatial extents of the 8 “master key filters” using Method C2 in Section 3.4, based on minimizing the discrete $l_{1}$-norm between the idealized receptive field models and the normalized versions of the learned filters according to (120). (Idealized 1): Non-centered vertical difference operator $(\delta_{y+}T)(m,n;\;\sigma_{1},\sigma_{1})$. (Idealized 2): Non-centered horizontal difference operator $(\delta_{x-}T)(m,n;\;\sigma_{2},\sigma_{2})$. (Idealized 3): Non-centered vertical difference operator $(\delta_{y-}T)(m,n;\;\sigma_{3},\sigma_{3})$. (Idealized 4): Non-centered horizontal difference operator $(\delta_{x+}T)(m,n;\;\sigma_{4},\sigma_{4})$. (Idealized 5): Centered horizontal difference operator $(\delta_{x}T)(m,n;\;\sigma_{5},\sigma_{5})$. (Idealized 6): Centered vertical difference operator $(\delta_{y}T)(m,n;\;\sigma_{6},\sigma_{6})$. (Idealized 7): Sharpening operation $1-\gamma_{7}\,(\nabla_{5}^{2}T)(m,n;\;\sigma_{7},\sigma_{7})$. (Idealized 8): Pure spatial smoothing $T(m,n;\;\sigma_{8},\sigma_{8})$. (The numerical values of $\sigma_{i}$ for $i\in\{1,2,3,4,5,6,8\}$ are according to Table 10, and with $\sigma_{7}$ additionally according to Table 13.) (Note that the contrasts of Filters 2 and 3 are reversed in relation to the sign conventions for corresponding “master key filters”.)
3.4 Method C2: $l_{1}$-norm-based fitting to idealized model for scale-space operations with the same scale parameter along the horizontal and the vertical directions
As a complement to the above analysis, let us also consider the special case when scale parameters are required to be the same in the horizontal and the vertical directions $\sigma_{x,i}=\sigma_{y,i}=\sigma_{i}$, for which we then determine each scale estimate $\hat{\sigma}_{i}$ from the criterion
$\hat{\sigma}_{i}=\operatorname{argmin}_{\hat{\sigma}_{i}}\|\Delta h_{i,\scriptsize\mbox{norm}}(\cdot,\cdot;\;\sigma_{i},\sigma_{i})\|_{1}=\\ =\operatorname{argmin}_{\sigma_{i}}\|h_{i,\scriptsize\mbox{ideal}}(\cdot,\cdot;\;\sigma_{i},\sigma_{i})-h_{i,\scriptsize\mbox{norm}}(\cdot,\cdot)\|_{1}.$ (120)
Table 10 shows the result of computing scale estimates for the idealized receptive field models in this way. Figure 9 shows visualizations of the resulting receptive fields.
Idealized 1(D1) Idealized 2(D1) Idealized 3(D1) Idealized 4(D2)
Refer to caption Refer to caption Refer to caption Refer to caption
Idealized 5(D1) Idealized 6(D1) Idealized 7(D2) Idealized 8(D1)
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 10: Idealized models of the 8 “master key filters” in Figure 2, in terms of combinations pure difference operators with spatial smoothing filters based on the discrete analogue of the Gaussian kernel, with the amount of spatial smoothing determined from the spatial extents of the 8 “master key filters” using Method D1 in Section 3.5, based on minimizing the discrete $l_{2}$-norm between the idealized receptive field models and the normalized versions of the learned filters and using different values of the scale parameters in the horizontal and the vertical directions according to (121). (Idealized 1): Non-centered vertical difference operator $(\delta_{y+}T)(m,n;\;\sigma_{x,1},\sigma_{x,1})$. (Idealized 2): Non-centered horizontal difference operator $(\delta_{x-}T)(m,n;\;\sigma_{x,2},\sigma_{x,2})$. (Idealized 3): Non-centered vertical difference operator $(\delta_{y-}T)(m,n;\;\sigma_{x,3},\sigma_{x,3})$. (Idealized 4): Non-centered horizontal difference operator $(\delta_{x+}T)(m,n;\;\sigma_{x,4},\sigma_{x,4})$. (Idealized 5): Centered horizontal difference operator $(\delta_{x}T)(m,n;\;\sigma_{x,5},\sigma_{x,5})$. (Idealized 6): Centered vertical difference operator $(\delta_{y}T)(m,n;\;\sigma_{x,6},\sigma_{x,6})$. (Idealized 7): Sharpening operation $1-\gamma_{7}\,(\nabla_{5}^{2}T)(m,n;\;\sigma_{7},\sigma_{7})$. (Idealized 8): Pure spatial smoothing $T(m,n;\;\sigma_{x,8},\sigma_{x,8})$. (The numerical values of $\sigma_{i}$ for $i\in\{1,2,3,4,5,6,8\}$ are according to Table 11, and with $\sigma_{7}$ additionally according to Table 13.) (Note that the contrasts of Filters 2 and 3 are reversed in relation to the sign conventions for corresponding “master key filters”.)
Filter $\hat{\sigma}_{x,i}$ $\hat{\sigma}_{y,i}$
1 0.491 0.722
2 0.581 0.519
3 0.483 0.503
4 0.500 0.000
5 1.300 1.004
6 0.984 1.074
8 0.615 0.608
Table 11: Estimates of the horizontal and vertical scale parameters $\hat{\sigma}_{x,i}$ and $\hat{\sigma}_{y,i}$ computed according to Method D1, based on minimizing the discrete $l_{2}$-norm between the idealized receptive field model $h_{i,\scriptsize\mbox{ideal}}(\cdot,\cdot;\;\sigma_{x,1},\sigma_{y,1})$ according to (93)–(100) and the normalized learned filter $h_{i,\scriptsize\mbox{norm}}(\cdot,\cdot)$ according to (63)–(68) and (73).
3.5 Method D1: $l_{2}$-norm-based fitting to idealized model for scale-space operations with different scale parameters along the horizontal and the vertical directions
In analogy with the above $l_{1}$-norm-based method in Section 3.3, for determining separate scale parameters in the horizontal and the vertical directions for idealized receptive field models, we can also an $l_{2}$-norm-based method from the criterion
$(\hat{\sigma}_{x,i},\hat{\sigma}_{y,i})=\operatorname{argmin}_{(\hat{\sigma}_{x,i},\hat{\sigma}_{y,i})}\|\Delta h_{i,\scriptsize\mbox{norm}}(\cdot,\cdot;\;\sigma_{x,i},\sigma_{y,i})\|_{2}=\\ =\operatorname{argmin}_{(\hat{\sigma}_{x,i},\hat{\sigma}_{y,i})}\|h_{i,\scriptsize\mbox{ideal}}(\cdot,\cdot;\;\sigma_{x,i},\sigma_{y,i})-h_{i,\scriptsize\mbox{norm}}(\cdot,\cdot)\|_{2}$ (121)
with the idealized receptive field models $h_{i,\scriptsize\mbox{ideal}}(\cdot,\cdot;\;\sigma_{x,1},\sigma_{y,1})$ again according to (93)–(100) and the normalized “master key filters” $h_{i,\scriptsize\mbox{norm}}(\cdot,\cdot)$ according to (63)–(68) and (73).
Table 11 shows the horizontal and the vertical scale estimates obtained in this way. Figure 10 shows visualizations of the corresponding idealized receptive fields.
Filter $\hat{\sigma}_{i}$
1 0.644
2 0.558
3 0.495
4 0.380
5 1.193
6 1.038
8 0.612
Table 12: Estimates of the joint horizontal and vertical scale parameters $\hat{\sigma}_{i}$ computed according to Method D2, based on minimizing the discrete $l_{2}$-norm between the idealized receptive field model $h_{i,\scriptsize\mbox{ideal}}(\cdot,\cdot;\;\sigma_{x,1},\sigma_{y,1})$ according to (93)–(100) and the normalized learned filter $h_{i,\scriptsize\mbox{norm}}(\cdot,\cdot)$ according to (63)–(68) and (73), under the assumption that the values of the horizontal and the vertical scale parameters are the same $\sigma_{x,i}=\sigma_{y,i}=\sigma_{i}$.
Idealized 1(D2) Idealized 2(D2) Idealized 3(D2) Idealized 4(D2)
Refer to caption Refer to caption Refer to caption Refer to caption
Idealized 5(D2) Idealized 6(D2) Idealized 7(D2) Idealized 8(D2)
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 11: Idealized models of the 8 “master key filters” in Figure 2, in terms of combinations pure difference operators with spatial smoothing filters based on the discrete analogue of the Gaussian kernel, with the amount of spatial smoothing determined from the spatial extents of the 8 “master key filters” using Method D2 in Section 3.6, based on minimizing the discrete $l_{2}$-norm between the idealized receptive field models and the normalized versions of the learned filters and using the same values of the scale parameters in the horizontal and the vertical directions according to (120). (Idealized 1): Non-centered vertical difference operator $(\delta_{y+}T)(m,n;\;\sigma_{1},\sigma_{1})$. (Idealized 2): Non-centered horizontal difference operator $(\delta_{x-}T)(m,n;\;\sigma_{2},\sigma_{2})$. (Idealized 3): Non-centered vertical difference operator $(\delta_{y-}T)(m,n;\;\sigma_{3},\sigma_{3})$. (Idealized 4): Non-centered horizontal difference operator $(\delta_{x+}T)(m,n;\;\sigma_{4},\sigma_{4})$. (Idealized 5): Centered horizontal difference operator $(\delta_{x}T)(m,n;\;\sigma_{5},\sigma_{5})$. (Idealized 6): Centered vertical difference operator $(\delta_{y}T)(m,n;\;\sigma_{6},\sigma_{6})$. (Idealized 7): Sharpening operation $1-\gamma_{7}\,(\nabla_{5}^{2}T)(m,n;\;\sigma_{7},\sigma_{7})$. (Idealized 8): Pure spatial smoothing $T(m,n;\;\sigma_{8},\sigma_{8})$. (The numerical values of $\sigma_{i}$ for $i\in\{1,2,3,4,5,6,8\}$ are according to Table 12, and with $\sigma_{7}$ additionally according to Table 13.) (Note that the contrasts of Filters 2 and 3 are reversed in relation to the sign conventions for corresponding “master key filters”.)
3.6 Method D2: $l_{2}$-norm-based fitting to idealized model for scale-space operations with the same scale parameter along the horizontal and the vertical directions
As a complement to the above analysis, let us also consider the special case when scale parameters are required to be the same in the horizontal and the vertical directions $\sigma_{x,i}=\sigma_{y,i}=\sigma_{i}$, for which we then determine each scale estimate $\hat{\sigma}_{i}$ from the criterion
$\hat{\sigma}_{i}=\operatorname{argmin}_{\hat{\sigma}_{i}}\|\Delta h_{i,\scriptsize\mbox{norm}}(\cdot,\cdot;\;\sigma_{i},\sigma_{i})\|_{2}=\\ =\operatorname{argmin}_{\sigma_{i}}\|h_{i,\scriptsize\mbox{ideal}}(\cdot,\cdot;\;\sigma_{i},\sigma_{i})-h_{i,\scriptsize\mbox{norm}}(\cdot,\cdot)\|_{2}.$ (122)
Table 12 shows the result of computing scale estimates for the idealized receptive field models in this way. Figure 11 shows visualizations of the resulting idealized receptive fields.
Norm $\hat{\sigma}_{7}$ $\hat{\gamma}_{7}$
$l_{1}$ 0.654 0.522
$l_{2}$ 0.675 0.526
Table 13: Estimates of the parameters $\sigma_{7}$ and $\gamma_{7}$ in the sharpening filter model (99) of Filter 7, computed with respect to minimizing either the discrete $l_{1}$-norm or the discrete $l_{2}$-norm of the difference between the idealized receptive field model and the normalized filter (72) according to (123) and (124).
3.7 Two-parameter modelling of Filter 7 as a sharpening filter
For Filter 7, for which the idealized model in (99) comprises a complementary scaling parameter $\gamma_{7}$ in addition to the scale parameter $\sigma_{7}$, we determine these two parameters simultaneously, by joint minimization over either the discrete $l_{1}$-norm or the discrete $l_{2}$-norm according to
$(\hat{\sigma}_{7},\hat{\alpha}_{7})=\operatorname{argmin}_{(\sigma_{7},\alpha_{7})}\|\Delta h_{7,\scriptsize\mbox{norm}}(\cdot,\cdot;\;\sigma_{7},\alpha_{7})\|_{1}=\\ =\operatorname{argmin}_{(\sigma_{7},\alpha_{7})}\|h_{7,\scriptsize\mbox{ideal}}(\cdot,\cdot;\;\sigma_{7},\alpha_{7})-h_{i7\scriptsize\mbox{norm}}(\cdot,\cdot)\|_{1},$ (123)
$(\hat{\sigma}_{7},\hat{\alpha}_{7})=\operatorname{argmin}_{(\sigma_{7},\alpha_{7})}\|\Delta h_{7,\scriptsize\mbox{norm}}(\cdot,\cdot;\;\sigma_{7},\alpha_{7})\|_{2}=\\ =\operatorname{argmin}_{(\sigma_{7},\alpha_{7})}\|h_{7,\scriptsize\mbox{ideal}}(\cdot,\cdot;\;\sigma_{7},\alpha_{7})-h_{i7\scriptsize\mbox{norm}}(\cdot,\cdot)\|_{2}.$ (124)
These methods are in the graphical illustrations above also referred to as “Method C2” and “Method D2”, respectively.
Table 13 shows numerical values of the resulting parameter estimates $\hat{\sigma}_{7}$ and $\hat{\gamma}_{7}$ obtained in this way.
Visualizations of the corresponding filter maps are given among the visualizations of the other filter maps for Method C2 and Method D2 in Figures 9 and 11, respectively.
Method A Method B Method C1 Method C2 Method D1 Method D2
Filter $\hat{\sigma}_{x,i}$ $\hat{\sigma}_{y,i}$ $\hat{\sigma}_{x,i}$ $\hat{\sigma}_{y,i}$ $\hat{\sigma}_{x,i}$ $\hat{\sigma}_{y,i}$ $\hat{\sigma}_{i}$ $\hat{\sigma}_{x,i}$ $\hat{\sigma}_{y,i}$ $\hat{\sigma}_{i}$
1 0.580 0.558 0.644 0.583 0.360 0.510 0.458 0.491 0.722 0.644
2 0.559 0.580 0.586 0.644 0.555 0.453 0.448 0.581 0.519 0.558
3 0.617 0.601 0.690 0.674 0.701 0.655 0.671 0.483 0.503 0.495
4 0.642 0.424 0.756 0.460 0.563 0.384 0.420 0.500 0.000 0.380
5 1.075 0.800 1.107 0.945 1.309 0.875 1.387 1.300 1.004 1.193
6 0.771 1.003 0.900 0.889 0.973 1.171 1.090 0.984 1.074 1.038
7 0.654 0.675
8 0.552 0.545 0.609 0.601 0.637 0.587 0.611 0.615 0.608 0.612
Table 14: Summary of the scale estimates obtained for the 8 “master key filters” using the different criteria using for computing them in different ways using: (i) Method A based on direct lookup of anisotropic scale estimates from variance-based spatial spread measures, (ii) Method B based on anisotropic matching of discrete spatial spread measures between the idealized receptive field models and the learned filters, (iii) Method C1 based on anisotropic discrete $l_{1}$-norm minimization, (iv) Method C2 based on isotropic discrete $l_{1}$-norm minimization, (v) Method D1 based on anisotropic discrete $l_{2}$-norm minimization and (vi) Method D2 based on isotropic discrete $l_{1}$-norm minimization. These methods have been explored for all the learned filters except for Filter 7, for which instead the two-parameter modelling scheme in Section 3.7, which also determines a scaling constant for the negative Laplacian-of-the-Gaussian filter in the sharpening operation.
Table 14 gives a summarizing overview of the scale estimates obtained from the different methods applied to the individual learned filters.
4 Interpretation of the analysis and modelling results in terms of scale-space theory for spatial receptive fields
4.1 The spaces spanned by linear combinations of the receptive field responses from the set of “master key filters”
A notable property of the learned receptive fields is that their orientational preferences are very closely aligned to the directions of Cartesian coordinate directions. In our preliminary investigations in Section 2.2, we noted that the unweighted spatial covariance matrices of the absolute values of the learned filter kernels were very close to diagonal (see Table 1). In our follow-up analysis in Section 2.5, we also noted that the analogous weighted spatial covariance matrices of the absolute values of the learned filter kernels were also very close to diagonal (see Table 6). This overall property then motivated us to state separable idealized models of the learned receptive fields according to (93)–(100), which we then in Section 3 demonstrated to provide qualitatively reasonable models of the shapes of the learned filters.
Given that natural image structures usually show a substantial variability with regard to orientations in the image domain, one may then ask how such a variability could be spanned by the deep network in which the learned filters constitute the primary computational primitives. From the property of the first-order directional derivative operator $\partial_{\varphi}$ in the direction $\varphi$
$\partial_{\varphi}=\cos\varphi\,\partial_{x}+\sin\varphi\,\partial_{y},$ (125)
it does, however, follow that the computation of derivatives in the horizontal and vertical directions, as performed by the operators $\partial_{x}$ and $\partial_{y}$, can be linear combinations span the space of all directions $\varphi$ in the 2-D image plane. Thus, since shapes of Filters 1–6 can be qualitatively well described by first-order Gaussian derivative kernels in the Cartesian coordinate directions, we could regard the depthwise combinations of the responses from the learned filters as having the potential ability to span a reasonable approximation of the space of the first-order directional derivatives in any direction $\varphi$, at two separate scales: both (i) the joint scale of Filters 1–4 and (ii) the joint scale of Filters 5–6.
Secondly, one may ask why there is a clear preference to first-order spatial derivatives in the learned “master key filters”, and not regarding e.g. second-order spatial derivatives, and why the receptive fields of Filters 1–4 are not fully centered, but instead shifted by approximately half a grid spacing along its preferred orientation in the image plane. From properties of the non-centered primitive first-order difference operators $\delta_{x+}$, $\delta_{x-}$, $\delta_{y+}$ and $\delta_{y-}$, we can, however, first of all by linear combinations construct their corresponding centered first-order difference operators according to
$\begin{split}\delta_{x}=\frac{\delta_{x+}+\delta_{x-}}{2},\end{split}$ (126)
$\begin{split}\delta_{y}=\frac{\delta_{y+}+\delta_{y-}}{2}.\end{split}$ (127)
We can also construct the corresponding centered second-order difference operators along the coordinate directions as
$\begin{split}\delta_{xx}=\delta_{x+}-\delta_{x-},\end{split}$ (128)
$\begin{split}\delta_{yy}=\delta_{y+}-\delta_{y-}.\end{split}$ (129)
Thus, by linear combinations of the output from Filters 1–4, it should be possibly to to a certain degree of approximation span the space of second-order directional derivatives along the Cartesian coordinate directions at the joint scale of Filters 1–4.
If we would like to span the space of second-order directional derivatives in any direction $\varphi$ according to
$\partial_{\varphi\varphi}=\cos^{2}\varphi\,\partial_{xx}+2\cos\varphi\,\sin\varphi\,\partial_{xy}+\sin^{2}\varphi\,\partial_{yy},$ (130)
we would, however, need to have access to a discrete approximation of the mixed second-order derivative operator $\partial_{xy}$, in order to be able to compute a reasonable discrete approximation of the second-order directional derivative in any direction in the image plane.
Given this theoretical motivation, one could thus raise the question of whether the performance of depthwise-separable deep networks with the receptive field shapes replaced by idealized receptive field models could be improved by extending the set of “master key filters” with a filter that constitutes a discrete approximation of the mixed second-order Gaussian derivative operator at the joint scale of Filters 1–4:
$\begin{split}h_{9,\scriptsize\mbox{ideal}}(m,n;\;\sigma_{x,9},\sigma_{y,9})=(\delta_{xy}T)(m,n;\;\sigma_{x,9},\sigma_{y,9}).\end{split}$ (131)
Concerning these interpretations, it should finally be emphasized that in depthwise-separable CNN:s, no direct linear combinations of the depthwise filters are performed. The depthwise convolutions are applied on distinct input feature maps, and the following pointwise layers linearly combine the outputs of these convolutions. Hence, the interpretations of the receptive field spanning different subspaces should be interpreted in terms of gross properties at the level of the population of the receptive fields only, and not concerning individual receptive fields.
4.2 Implications when designing Gaussian derivative networks
In previous studies of Gaussian derivative networks, the receptive fields in these networks have mainly been formulated as linear combinations of spatial derivatives up to orders 2 or 3, and with the underlying Gaussian derivative responses computed at the same spatial scale level, see Jacobsen et al. (2016), Lindeberg (2021b, 2022), Pintea et al. (2021), Sangalli et al. (2022), Penaud-Polge et al. (2022), Gavilima-Pilataxi and Ibarra-Fiallo (2023) and Perzanowski and Lindeberg (2025). Given the results from the presented analysis of the set of “master key filters”, one may, however, raise the question of whether the performance of Gaussian derivative networks could be improved by:
  • Combining receptive field responses from two scale levels instead of one scale level in each Gaussian derivative layer, as motivated by the qualitatively different scale levels of the groups of filters with indices $\{1,2,3,4,8\}$ and $\{5,6\}$, respectively.
  • Adding a zero-order term to those networks that have not previously made use of such a term, as motivated by the qualitative shape of Filter 8 in the set of “master key filters”.
  • Adding an isotropic sharpening term, as motivated by the qualitative shape of Filter 7 in the set of “master key filters”.
In this context, it should be specifically noted that the use of linear combinations of Gaussian derivative responses in the layers of Gaussian derivative network constitutes a computational operation with large structural similarity to the combination of a set of basic filtering primitives in the depthwise combination of layers in depthwise-separable networks.
4.3 Implications when using idealized models of “master key filters” in depthwise-separable deep networks
As previously discussed in Section 4.1, it could be interesting to investigate if the addition of an additional filter, corresponding to a discrete approximation of the mixed second-order derivative $\partial_{xy}$ at the joint scale of Filters 1–4, could be advantageous to depthwise-separable network, by making it theoretically possible also to a certain degree of approximation span the space of second-order directional derivative operators according to (130), by in this way enabling formal capture of a significantly richer set of spatial image structures.
As can be seen from the results in Section 3 of estimating the scale values of idealized filter models to approximate the 8 “master key filters” by discrete scale-space operations, the numerical values of the resulting scale estimates can, however, vary notable depending on what criteria are used for computing the scale estimates. For Filter 8, the variability depending on the different criteria is marginal. For Filters 1–6, the variability is, however, significant, because of notable deviations between the idealized receptive field models and the learned filters.
One may hence ask what are the possible predictive properties of these different criteria with regard to the application of replacing the learned filter models by idealized receptive fields in depthwise-separable deep networks. Ultimately, the validity of these idealized approximations is determined by the performance in actual experiments.
A natural next step to consider would also be to use the proposed structural forms of idealized receptive fields according to the presented methods as initial seeds, to then optimize the choice of parameter settings for the idealized filters in a complementary training stage, in a corresponding way as learning of the scale levels (Pintea et al. (2021, Penaud-Polge et al. (2022, Saldanha et al. 2021, Yang et al. 2023, Basting et al. 2024, Perzanowski and Lindeberg (2025) can be applied to regular Gaussian derivative networks.
5 Experimental results
To validate the predictive properties of our different idealized filter modelling approaches, we conducted a series of experiments on the ImageNet dataset. Our experimental design follows a similar methodology as described in Babaiee et al. (2025b), where we systematically replace the learned depthwise filters in DS-CNN architectures with idealized approximations derived from our theoretical analysis.
5.1 Comparative evaluation of idealized filter modelling methods
We begin by evaluating the predictive performance of all the six modelling methods (Methods A, B, C1, C2, D1, and D2) described in Section 3. For each method, we replaced all the depthwise filters in the ConvNeXt v2 Tiny architecture with linear approximations of the form $af^{\prime}+b$, where $f^{\prime}$ represents the closest match among our 8 idealized filters, following the procedure outlined in Babaiee et al. (2025b)
The linear approximation parameters $a$ and $b$ were determined by minimizing the Euclidean distance between the original learned filter $f_{c}$ and the transformed idealized filter $af^{\prime}_{c}+b$, using the closed-form solution presented in Equations (1)-(3) in the original work. This approach ensures that each learned filter is optimally approximated by a scaled and shifted version of one of our 8 fundamental filter types.
Table 15 presents the ImageNet Top-1 accuracy results for the ConvNeXt v2 Tiny model when its depthwise filters are replaced with idealized approximations from each of our six methods, without any subsequent fine-tuning of the rest of the parameters.
Method A B C1 C2 D1 D2
Top-1 (%) 63.958 65.700 62.697 60.972 62.330 63.804
Table 15: ImageNet Top-1 accuracy comparison for different idealized filter modelling methods applied to ConvNeXt v2 Tiny without fine-tuning.
The results demonstrate that Method B, based on matching discrete weighted variance-based spatial spread measures between idealized receptive field models and learned filters, achieves the highest predictive accuracy at 65.70%. This finding is particularly significant, as it suggests that the discrete modelling approach, which accounts for the inherent discretization effects in the learned filters, provides superior approximation quality compared to methods based on continuous Gaussian derivative models or direct norm minimization.
5.2 Validation through frozen filter training
Based on the comparative results, we selected Method B for further investigation of its practical applicability in depthwise-separable networks. We conducted experiments where the ConvNeXt v2 Tiny model was initialized with our 8 idealized filters (determined using Method B parameter estimates from Table 8) and trained from scratch on ImageNet with the depthwise filters remaining frozen throughout the 300-epoch training process.
For determining the filter type assignment at initialization, we analyzed the fully trained ConvNeXt v2 Tiny model and classified each depthwise filter according to its closest match among our 8 idealized filter types using the linear approximation procedure described before. These filter type assignments were then fixed throughout the subsequent training process. All the other network parameters were initialized from the FCMAE (fully convolutional masked autoencoder) pre-trained ConvNeXt v2 Tiny weights, following the same initialization procedure as described in the original ConvNeXt v2 paper. For comparative purposes, we also trained a standard ConvNeXt v2 Tiny model using the FCMAE pretrained initialization, following the exact hyperparameters and training procedures from the original code repository. This baseline model achieved 82.7% accuracy and serves as our reference for evaluating the performance impact of our idealized filter constraints.
Table 16 compares the performance of different initialization strategies for the ConvNeXt v2 Tiny architecture.
Configuration Top-1 Accuracy (%)
Original ConvNeXt v2 Tiny 82.7
Frozen 8 master key filters 82.7
Frozen 8 filters from Method B 82.5
With learning of the scale parameters 82.6
Table 16: ImageNet Top-1 accuracy comparison for ConvNeXt v2 Tiny with different filter initialization strategies.
The results reveal a remarkable consistency across different initialization approaches. The model initialized with our Method B idealized filters achieves 82.5% accuracy, representing only a 0.2% decrease compared to the original architecture. This minimal performance degradation, despite using only 8 distinct filter types across all the depthwise layers, provides strong empirical support for our theoretical analysis and demonstrates the effectiveness of our scale-space-based modelling approach.
Notably, the frozen master key filters from Babaiee et al. (2025b) achieve identical performance to the fully trainable baseline, while our theoretically-derived approximations maintain nearly equivalent results. This suggests that our idealized receptive field models successfully capture the essential computational properties of the learned filters.
5.3 Learning of the scale parameters
To investigate whether the theoretical constraints imposed by our idealized models limit the representational capacity of the network, we conducted an additional experiment where the 8 filter shapes from Method B were frozen, but their scale parameters $\sigma_{x,i}$ and $\sigma_{y,i}$ were made trainable. This approach maintains the theoretical structure of our scale-space operators, while allowing adaptation to the specific requirements of the layers and architecture.
The model with learnable scale parameters achieved 82.6% accuracy, representing a slight improvement of 0.1% over the fully frozen configuration. This result suggests that while our theoretical parameter estimates provide reasonable initial approximations, gradient-based optimization can yield marginal performance gains.
To understand how the network adapts the scale parameters during training, we analyzed the learned scale parameter distributions across all depthwise layers. Table 17 presents aggregated statistics for the scale parameters $\sigma_{x}$ and $\sigma_{y}$ for each of the 8 filter types after training completion.
Filter NN $\sigma_{x}$ mean $\sigma_{x}$ std $\sigma_{y}$ mean $\sigma_{y}$ std
1 164 0.505 0.189 0.538 0.254
2 132 0.535 0.234 0.538 0.319
3 211 0.523 0.195 0.548 0.222
4 167 0.534 0.242 0.549 0.307
5 154 0.625 0.098 0.677 0.176
6 705 0.657 0.132 0.623 0.085
7 915 0.588 0.274 0.589 0.197
8 4176 0.781 0.498 0.818 0.486
Table 17: Statistics of learned scale parameters across all the layers for the 8 filter types in terms of the means and the standard deviations over the sets of the learned filters. NN indicates the number of total filter instances using each type.
A pattern emerges from this analysis: for most filter types (Filters 1-7), the standard deviation of the learned scale parameters is relatively low, indicating that the network converges to similar scale parameter values across different instances of each filter type. This suggests that these computational primitives—corresponding to various derivative operations and sharpening filters—have fairly constrained optimal parameter ranges.
In contrast, Filter 8, which corresponds to the pure discrete Gaussian smoothing kernel $T(m,n;\sigma_{x,8},\sigma_{y,8})$, exhibits substantially higher standard deviations (0.498 and 0.486) while also being the most frequently used filter type (4176 instances). This high variance indicates that the network benefits from having different spatial smoothing rates for the Gaussian smoothing operation across different filter instances.
Notably, despite these variations in the learned scale parameters—particularly the deviation from our theoretical estimates and the high variance in Filter 8—the overall network accuracy remained essentially unchanged (82.6% vs 82.5%). This suggests that while the specific scale parameters may vary, the fundamental computational structure captured by our 8 idealized filter types is sufficient to maintain performance, indicating that the filter shapes themselves are more critical than their precise parameterization.
6 Summary and discussion
We have presented an in-depth analysis of different ways of modelling the 8 “master key filters” extracted from the ConvNeXt deep learning architecture by Babaiee et al. (2025a). For this purpose, we have first computed quantitative measures of the spatial extent of the learned filters in terms of both unweighted and weighted spatial spread measures, giving support for modelling the receptive fields of these filters with separable filtering operations along the Cartesian coordinate directions, and using non-centered difference operators for modelling Filters 1–4 and centered difference operators for modelling Filters 5–6, in combination with a spatial smoothing operations with possibly different values of the scale parameters along the different coordinate directions, and possibly different values of the scale parameters for the individual filters.
The presented modelling and analysis results show that the predictive properties of the scale-space framework to derive canonical models for spatial receptive fields in terms of Gaussian smoothing and Gaussian derivatives do qualitatively well generalize from the previously axiomatic formulations regarding the first layer of visual processing, to the qualitative shapes of the receptive fields learned in certain families of depthwise-separable deep networks. Conceptually, these results are also compatible with previous applications of the use Gaussian derivative operations as the computational primitives in deep networks, as developed by Jacobsen et al. (2016), Lindeberg (2021b, 2022), Pintea et al. (2021), Sangalli et al. (2022), Penaud-Polge et al. (2022), Gavilima-Pilataxi and Ibarra-Fiallo (2023) and Perzanowski and Lindeberg (2025).
To explore the influence of different choices for formulating mathematical criteria to model the learned receptive fields by idealized scale-space operations, we have specifically explored four main paths of such methods:
  • Method A based on measuring the spatial extent of the learned filters by weighted variance-based spatial spread measures and then predicting the values of the scale parameters based on properties of continuous Gaussian derivatives.
  • Method B based on matching discrete weighted variance-based spatial spread measures between the idealized receptive field models and the learned filters.
  • Methods C1 and C2 based on minimizing the discrete $l_{1}$-norm between the idealized receptive field models and the learned filters.
  • Methods D1 and D2 based on minimizing the discrete $l_{2}$-norm between the idealized receptive field models and the learned filters.
As we have found from the resulting experiments, for some of the learned filters, the scale estimates may vary notable depending on the choice of matching criterion, while for Filter 8 the difference is minor. The main reason for this variability in discrepancy is because of a varying degree of agreement between the idealized models and the learned filters. For judging which of these methods has the best predictive properties with regard to the performance, when used for replacing the computational primitives in depthwise-separable deep networks, this performance should therefore be evaluated experimentally. Specifically, to be less dependent on the choice of mathematical criterion for model fitting, we have proposed to use the presented theory for formulating parameterized primitives for the idealized receptive field models, and then learning these parameters by deep learning, in a corresponding way as learning of the scale levels is used in Gaussian derivative networks.
The experimental results of replacing the 8 “master key filters” in the ConvNeXt v2 Tiny deep network by idealized discrete scale-space filters obtained in this way show that almost as good results are obtained by using the idealized discrete scale-space filters as when using either the 8 “master key filters” or the filters learned by back propagation in the original deep network architecture. Thereby, the results demonstrate that the filters learned in a reduced version of the modern and state-of-the-art deep network architecture ConvNeXt appear to be very similar to the filters obtained by axiomatically determined scale-space theory. Furthermore, we have shown that the additional increase in performance obtained by learning the scale parameters is marginal for the ConvNeXt v2 Tiny model, thus demonstrating the very good predictive ability of Method B when applied to the filters learned in this deep learning architecture.
Notably the amount of spatial smoothing is very small in the idealized models of the 8 “master key filters”. In this modelling work, we have throughout modelled this amount of smoothing with spatial filters based on the discrete analogue of the Gaussian kernel (Lindeberg 1990) combined with small-support difference operators, motivated by the theoretical results in Lindeberg (2024, 2025), that demonstrate that this discretization method is the discretization method that best preserve a transfer of scale-space properties from a continuous to a discrete image domain, and the experimental results in Perzanowski and Lindeberg (2025), that show that this discretization method overall gives the best performance among 5 different discretization methods, when approximating Gaussian derivative operators for application in Gaussian derivative networks.
Concerning the weighted variance-based spread measures for estimating the spatial extents of the kernels, we have in this work throughout used fixed values of the scale parameters for the spatial weighting functions, coarsely determined from the spatial extents of the learned filters as observed by visual inspection. A more general approach in this respect could instead be to vary these scale parameters, or to try to adapt these scale parameters in an iterative manner.
Another conceptual simplification that we have done in the above model fitting schemes is that we have normalized the learned receptive fields that are approximated by first-order derivatives in the horizontal xx–direction to having the same response to the monomial $x$ as a corresponding first-order derivative derivative $g_{x}(x,y;\,\sigma)$. Similarly, we have a pre-processing step to the model fitting stage normalized the learned receptive fields that are approximated by first-order derivatives in the vertical yy–direction to having the same response to the monomial $y$ as a corresponding first-order derivative derivative $g_{y}(x,y;\,\sigma)$. Furthermore, we determined offset constants for Filters 7–8 to minimize their variance-based spatial spread measures, and normalized their responses to a constant unit function to be equal to 1, to make the resulting normalized filters more similar to either a pure spatial smoothing operation or a local sharpening operation. In these ways, we reduced the dimensionality of the following model fitting search, although the resulting model fitting for the $l_{1}$- and $l_{2}$-norm-based modelling schemes may possibly not be as accurate as if it would be extended to also involve a search over the amplitudes of the receptive fields. A possible extension to future work could therefore involve performing such higher-dimensional optimizations over the spaces of the parameters of the idealized receptive fields.
Finally, let us remark that although the work in this paper has been focused on modelling and analyzing the 8 specific “master key filters” extracted by Babaiee et al. (2025a), we propose that a similar methodology could be used for modelling other sets of “master key filters” obtained from other deep networks as well as biological visual receptive fields recorded by neurophysiological measurements.